### v1

# Minimal Degrees, Volume Growth, and Curvature Decay on Complete Kähler Manifolds

Abstract.
We consider noncompact complete Kähler manifolds with nonnegative bisectional curvature. Our main results are: 1. Precise estimates among refined minimal degree of polynomial growth holomorphic functions and holomorphic volume forms, $\operatorname{AVR}$ (asymptotic volume ratio) and $\operatorname{ASCD}$ (average of scalar curvature decay) are established. 2. The Lyapunov asymptotic behavior of the Kähler-Ricci flow can be described in terms of polynomial growth holomorphic functions. This provides a unifying perspective that bridges the two distinct proofs of Yau’s uniformization conjecture. These resolve two conjectures made by Yang.
1. Introduction
As part of Yau’s program to study complex manifolds of parabolic type, he proposed the following well-known uniformization conjecture in 1970s:
Conjecture 1.1 (Yau’s Uniformization conjecture, [29]).
A complete noncompact Kähler manifold $(M^{n},g)$ with positive bisectional curvature is biholomorphic to $\mathbb{C}^{n}$.
In the maximal volume growth case, Liu’s breakthrough [16] confirmed the conjecture by combining Gromov-Hausdorff convergence techniques with the three-circle theorem developed in [13]. An alternative proof, based on the Kähler-Ricci flow, was later provided by Lee-Tam [8], building on results in Chau-Tam [3].
These advances led to a deeper understanding of the structure of such manifolds. Notably, through the works of [20, 21, 24, 11, 12, 17], the following conjecture of Ni has been established as a theorem:
Theorem 1.2 (Corollary 3.2 in [20], Theorem 1.2 in [24], Theorem 2 in [12], Theorem 1.4 in [11], Corollary 2.16 in [17]).
Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature. Assume that the universal cover of $M$ does not split. Then the following conditions are equivalent:
(1) $M$ is of maximal volume growth, i.e.
$\operatorname{AVR}(M,g)=\lim_{r\to\infty}\frac{\operatorname{Vol}\left(B(p,r)\right)}{\omega_{2n}r^{2n}}=\nu>0.$
Here $\omega_{2n}$ is the volume of the unit ball in $\mathbb{C}^{n}$.
(2) There exists a nonconstant polynomial growth holomorphic function, i.e. $\mathcal{O}_{P}(M)\neq\mathbb{C}$.
(3) The average scalar curvature decay is finite, i.e.
$\operatorname{ASCD}(M,g)=\limsup\limits_{r\to\infty}r^{2}\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p,r)}S$
is finite. Here $S$ is the scalar curvature.
Remark 1.3.
(1) Note that both $\operatorname{AVR}(M,g)$ and $\operatorname{ASCD}(M,g)$ are independent of the choice of $p$.
(2) A very recent result of Liu [18] implies that for any $(M,g)$ with nonnegative bisectional curvature, the $\limsup$” in the definition of $\operatorname{ASCD}(M,g)$ above can be replaced by $\lim$”, see also [22].
In his thesis [27], Yang defined:
Definition 1.4.
Let $(M^{n},g)$ be a complete Kähler manifold with nonnegative bisectional curvature, for a fixed $p\in M$, define
[…]
where $\mathcal{K}_{M}$ is the canonical line bundle on $M$, $\mathcal{O}_{P}(M)$ is the space of all polynomial growth holomorphic functions on $M$ and $H^{0}_{P}(M,\mathcal{K}_{M})$ is the space of all polynomial growth holomprhic $n$-forms on $M$. Set $d_{\operatorname{min}}=+\infty$ or $D_{\operatorname{min}}=+\infty$ if $M$ does not admit any nonconstant holomorphic functions of polynomial growth or $\mathcal{K}_{M}$ admits no nonzero holomorphic section of polynomial growth.
Now assume $(M^{n},\omega)$ satisfies the assumption in Theorem 1.2 and of maximal volume growth. Since the Kodaira dimension $K(M^{n})=n$, $\mathcal{O}_{P}(M)$ is “holomorphically regular” in the sense that we can always find “local coordinate by global polynomial growth functions”. The refined minimal degree can be also defined in [27] by
$\overrightarrow{d_{\operatorname{min}}}(p):=\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right):=\inf_{\{f_{1},\cdots,f_{n}\}}\left\{\limsup_{x\rightarrow\infty}\frac{\log|f_{1}(x)|}{\log d(x,p)},\cdots,\limsup_{x\rightarrow\infty}\frac{\log|f_{n}(x)|}{\log d(x,p)}\right\}$
where the infimum is taken among any $n$-tuple of global holomorphic functions that gives local coordinates at $p$ with the corresponding $\limsup_{x\rightarrow\infty}\frac{\log|f_{i}(x)|}{\log d(x,p)}$ arranged in a non-decreasing order for $1\leq i\leq n$, we denote this set by $\mathcal{O}_{P,p}(M)$. In other words, for any $1\leq k\leq n$ we have $d_{\operatorname{min}}^{(k)}=\inf_{f_{k}}\limsup_{x\to\infty}\frac{\log|f_{k}(x)|}{\log d(x,p)}$, where the infimum is taken among all possible $f_{k}$ that appears in the $k$-th component of some sequence in $\mathcal{O}_{P,p}(M)$. Note that apriori it’s unclear if $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$ can be obtained by an n-tuple holomorphic functions in $\mathcal{O}_{P,p}(M)$. But obviously $d_{\operatorname{min}}=d_{\operatorname{min}}^{(1)}$.
Then Yang proposed the following conjecture in [27] on the relation between the above quantities which can be understood as the quantitative version of Theorem 1.2.
Conjecture 1.5 (Conjecture 2.5.6, Conjecture 2.5.8 in [27]).
Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature. Assume the universal cover of $M$ does not split. Then
(1) $\overrightarrow{d_{\operatorname{min}}}(p)$ can be realized by a n-tuple holomorphic functions in $\mathcal{O}_{P,p}(M)$ and is independent of the choice of $p$.
(2) $\operatorname{AVR}(M,g)=\prod\limits_{i=1}^{n}\frac{1}{d_{\operatorname{min}}^{(i)}}$.
(3) $D_{\operatorname{min}}=\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n$.
(4) $\operatorname{ASCD}(M,g)=4nD_{\operatorname{min}}$.
In the case of nonmaximal volume growth, (1),(2) are true because $d_{\min}=+\infty$ from Theorem 1.2, and $\operatorname{ASCD}=+\infty$.
Assume $M$ is of maximal volume growth, in [27], Yang verified the conjecture is true for $U(n)$-invariant Kähler metrics on $\mathbb{C}^{n}$. Liu proved (1) implicitly in [16]. In fact, he showed that any n-tuple of polynomial growth functions which are algebraically independent and of minimal degrees can form a global coordinate. Then recently in [18], Liu proved (2) and derived the explicit formula $\operatorname{ASCD}(M,g)=4n(\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n)$ by passing the geometric quantities to the tangent cones and exploiting the metric Kähler cone structure of the limit space. In [28], Yang proved that (3) holds for gradient expanding Kähler-Ricci solitons with nonnegative Ricci curvature by using the Poincaré coordinate introduced in [1], see Theorem 1.8.
Therefore, the general proof of (3) remained the final step toward resolving Conjecture 1.5 in full generality.
In this paper, we prove that (3) of Conjecture 1.5 is true, thereby completely resolving the conjecture. Precisely, we proved
Theorem A.
Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature. Suppose the universal cover of $M$ does not split. Then
$D_{\operatorname{min}}=\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n.$
Now we move to the behavior of the Kähler-Ricci flow on complete noncompact Kähler manifolds where $(M,g)$ has nonnegative bisectional curvature.
$\frac{\partial g(t)}{\partial t}=-\operatorname{Ric}(g(t)),\quad g(0)=g.$ (1.1)
When the initial metric $(M,g)$ has bounded curvature and Euclidean volume growth, the long-time behavior of (1.1) has been described very clearly. Let us summarize these results in the following:
Theorem 1.6 (Theorem 1.2, Proposition 3.2 in [3], Theorem 6.1 in [4]).
Suppose that $(M^{n},J,g)$ is a complete noncompact Kähler manifold, where $g$ has bounded nonnegative bisectional curvature and Euclidean volume growth with $\operatorname{AVR}(M,g)=\nu>0$. Then the following conclusions hold:
(1) The solution to the Kähler-Ricci flow with the initial metric $g$ exists for all $t\in[0,+\infty)$ and it has nonnegative bisectional curvature for any $t\geq 0$. Moreover, there exists a constant $C(n,\nu)$ such that $\|\operatorname{Rm}(x,t)\|\leq\frac{C(n,\nu)}{t}$ for any $x\in M$ and $t\in(0,+\infty)$.
(2) For any point $x\in M$, let $\{\lambda_{1}(x,t),\cdots,\lambda_{n}(x,t)\}$ denote the eigenvalues of Ricci curvature $\operatorname{Ric}(x,t)$ with respect to $g(t)$ in the nondecreasing order. Then $t\lambda_{i}(x,t)$ is nondecreasing on $t>0$, hence $\mu_{i}(x):=\lim_{t\rightarrow+\infty}t\lambda_{i}(x,t)$ exists.
(3) If $\mu_{1}(x)<\mu_{2}(x)<\cdots<\mu_{l}(x)$ are the distinct limits in (2), where $l\leq n$. Then $V=T_{x}^{(1,0)}(M)$ can be decomposed orthogonally with respect to $g$ as $V_{1}\oplus\cdots\oplus V_{l}$ so that
if $v$ is a nonzero vector in $V_{i}$ for some $1\leq i\leq l$, and let $v(t)=\frac{v}{|v|_{g(t)}}$, then
$\lim\limits_{t\to\infty}t\operatorname{Ric}(v(t),\overline{v(t)})=\mu_{i}$
and thus
$\lim\limits_{t\to\infty}\log\frac{|v|^{2}_{g(t)}}{|v|^{2}_{g}}=-\mu_{i}.$
Moreover, both convergences are uniform over all $v\in V_{i}\backslash\{0\}$.
$\sum\limits_{i=1}^{l}-\mu_{i}(x)\operatorname{dim}_{\mathbb{C}}V_{i}=\lim\limits_{t\to\infty}\log\frac{\det(g_{i\bar{j}}(x,t))}{\det(g_{i\bar{j}}(x,0))}.$ (1.2)
(5) Fix any point $p\in M$. Given any $t_{k}\rightarrow+\infty$, define $g_{k}(t)=\frac{1}{t_{k}}g(t_{k}t)$. The pointed sequence $(M^{n},J,g_{k}(t),p)$ sub-sequentially converges to a gradient expanding Kähler-Ricci soliton $(N,J_{\infty},h(t),O)$ where $t\in(0,+\infty)$ in the following sense:
  1. (i)
    After picking a subsequence still denoted by $t_{k}$, there exists an increasing sequence of open subsets $O\in U_{k}$, which exhausts $N$ and a family of diffeomorphisms $F_{k}:U_{k}\rightarrow F_{k}(U_{k})\subset M$ with $F_{k}(O)=p$.
  2. (ii)
    As $t_{k}\rightarrow\infty$, the sequence $(U_{k},F_{k}^{*}J,F_{k}^{*}(\frac{1}{t_{k}}g(t_{k}t)),p)$ converges smoothly to another sequence of complete Kähler manifolds $(N,J_{\infty},h(t),O)$ uniformly on compact sets of $N\times(0,\infty)$.
  3. (iii)
    $(N,J_{\infty},h(t))$ has nonnegative bisectional curvature for any $t>0$, and there exists a real-valued function $f\in C^{\infty}(N)$ such that $(N,J_{\infty},h(1))$ satisfies the expanding soliton equation
    $\operatorname{Ric}_{i\bar{j}}(h(1))+h_{i\bar{j}}(1)-f_{i\bar{j}}=0,\quad f_{ij}=f_{\bar{i}\bar{j}}=0.$ (1.3)
    Moreover, $\nabla_{h(1)}f(O)=0$ and the eigenvalues of Ricci curvature of $h(1)$ at $O$, arranged in the non-decreasing order, equal $\mu_{i}(p)$ for $1\leq i\leq n$.
Indeed, in (3) in Theorem 1.6, $\mu_{i}$ play the role of Lyapunov exponents in a dynamical-system interpretation on asymptotics of (1.1), it basically says that the Ricci curvature can be “simultaneously diagonalized” near $t=+\infty$ in some sense. See Theorem 4.1 in [3] for a precise statement.
Then Yang proposed the following conjecture in [27]:
Conjecture 1.7 (Conjecture 2.5.16 in [27]).
Let $(M^{n},g)$ be a complete noncompact Kähler manifold with bounded nonnegative curvature and Euclidean volume growth. Let $g(t)$, $t\in[0,+\infty)$, be the complete solution Kähler-Ricci flow with initial metric $g$, then
(1) $\mu_{i}(p)$ is independent of the choice of $p$ for any $1\leq i\leq n$.
(2) $\mu_{i}=d_{\operatorname{min}}^{(i)}-1$ for any $1\leq i\leq n$.
(3) $\operatorname{ASCD}(M,g(t))$ is invariant along $g(t)$.
In [28], Yang proved (1) using Shi’s curvature estimates. He further proved (2) for expanding Kähler-Ricci solitons with nonnegative Ricci curvature. Precisely,
Theorem 1.8 (Theorem 1.4 in [28]).
Let $(N^{n},J,O,g,f)$ be a complete noncompact gradient Kähler-Ricci soliton with nonnegative Ricci curvature, normalized so that $f=R+|\nabla^{g}f|^{2}$. Let $\mu_{1}\leq\cdots\leq\mu_{n}$ be the eigenvalues of Ricci curvature at $O$. Then
$\overrightarrow{d_{\operatorname{min}}}(q)=\{\mu_{1}+1,\cdots,\mu_{n}+1\}\text{ for all }q\in N;$ (1.4)
and
$D_{\operatorname{min}}=\sum\limits_{i=1}^{n}\mu_{i}.$ (1.5)
If $(M,g)$ has unbounded curvature, in [8], Lee-Tam established the following long-time existence result for the Kähler-Ricci flow:
Theorem 1.9 (Corollary 1 in [21], Theorem 1.5 in [8]).
Suppose $(M^{n},g)$ is a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth, $\operatorname{AVR}(M,g)=\nu>0$. Then there exists $C=C(n,\nu)>0$ depending only on $n,\nu$ such that there exists a smooth Kähler-Ricci flow solution $g(t)$ on $[0,+\infty)$ such that
$\|\operatorname{Rm}\|(g(t))\leq\frac{C(n,\nu)}{t}$
on $M\times(0,+\infty)$.
(2) $g(t)$ has nonnegative bisectional curvature.
(3) $\operatorname{AVR}(g(t))=\operatorname{AVR}(g(0))=\nu$ on $M\times(0,+\infty)$.
We emphasize that the uniqueness of such Kähler-Ricci flow solution remains unknown in this setting. Consequently, though these $\mu_{i}$’s can be defined for $M$ with unbounded curvature, a priori, we cannot guarantee that they are independent of the particular flow solution g(t).
In this paper, we prove the Conjecture 1.7 holds in general.
Theorem B.
Conjecture 1.7 holds.
Since Conjecture 1.7 is true, then $\mu_{i}$’s are indeed independent of $g(t)$ and thus well-defined for manifolds with unbounded curvature. Therefore we can remove the boundedness condition in Conjecture 1.7.
Corollary 1.10.
Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative curvature and Euclidean volume growth. Let $g(t)$ $(t\in[0,+\infty)$ be as complete solution Kähler-Ricci flow with initial metric $g$ as in Theorem 1.9, then
(1) $\mu_{i}(p)$ is independent of the choice of $p$ for any $1\leq i\leq n$.
(2) $\mu_{i}=d_{\operatorname{min}}^{(i)}-1$ for any $1\leq i\leq n$.
(3) $\operatorname{ACSD}(M,g(t))$ is invariant along $g(t)$.
Note that by Claim 4.2 in [16], the degrees of functions in $\mathcal{O}_{P}(M)$ also can be “simultaneously diagonalized” along the blowdown sequence, for details see Theorem 2.4 (5) in Section 2. As a by-product of Corollary 1.10, we can describe the Lyapunov regularity in (3) of Theorem 1.6 using the polynomial ring $\mathcal{O}_{P}(M)$. In some sense, it bridges the two distinct proofs of Yau’s uniformization conjecture in the case of maximal volume growth.
Corollary 1.11.
In (3) of Theorem 1.6, the basis of $V_{i}$ can be taken as $\left\{\frac{\partial}{\partial f_{i1}}(x),\cdots,\frac{\partial}{\partial f_{im_{i}}}(x)\right\}$. Here $\operatorname{dim}V_{i}=m_{i}$ is the multiplicity of $\mu_{i}$, $\sum\limits_{i=1}^{l}m_{i}=n$, and
$f_{is}\in\mathcal{O}_{P}(M)\text{ with }\operatorname{deg}(f_{is})=d_{\operatorname{min}}^{(i)}$
for any $1\leq s\leq m_{i}$.
Moreover, $\{f_{11},\cdots,f_{1m_{1}},\cdots,f_{l1},\cdots,f_{lm_{l}}\}$ can serve as a biholomorphism from $M$ onto $\mathbb{C}^{n}$.
This paper is organized as follows. Section 2 is some basic preliminary results and some simple conclusions that will be used later. In section 3, we will prove Theorem B by noting that any tangent cone of $M$ also arises as a tangent cone of corresponding expanding Kähler-Ricci solitons in Theorem 1.6. Also we prove Corollary 1.11. Section 4 contains the proof of Theorem A.
Acknowledgements
I would like to express my sincere gratitude to my advisor, Professor Yihu Yang, for providing constant support. I would like to extend my thanks to Professor Gang Liu for many valuable discussions.
2. Preliminary results
2.1. Structure of Kähler manifolds with maximal volume growth and the tangent cones
We first recall the main result in [16], and prove some relevant facts that will be used later.
Theorem 2.1.
Let $(M^{n},g,p)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth. Let $\operatorname{AVR}(M,g)=\nu>0$. Then
(1) $M$ is biholomorphic to $\mathbb{C}^{n}$.
(2) There exists an n-tuple polynomial growth holomorphic functions $(f_{1},\cdots,f_{n})$ serving as a proper biholomorphism onto $\mathbb{C}^{n}$. These functions satisfy
$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p,1)}f_{i}\overline{f_{j}}=\delta_{ij},\quad f_{i}(p)=0\quad\text{ for any }i,j\in\{1,\cdots,n\}.$ (2.1)
Moreover, any element in $\mathcal{O}_{P}(M)$ is a polynomial of these functions, meaning that $\mathcal{O}_{P}(M)\cong\mathbb{C}[f_{1},\cdots,f_{n}]$.
(3) Any two such global coordinates (consisting of n polynomial growth holomorphic functions satisfying (2.1), serving as a proper biholomorphism onto $\mathbb{C}^{n}$, and generating $\mathcal{O}_{P}(M)$) differ by a constant orthogonal transformation.
Proof.
The proof of these conclusions is implicitly contained in [11, 16]. For reader’s convenience, we give a thorough explanation. (1) and (2) are contained in Section 4 in [16].
The properness is also contained in [11]. It was shown that there exists a constant $D=D(n,\nu)$ such that let $(g_{1},\cdots,g_{k})$ be a linearly independent basis of $\mathcal{O}_{P}(M)$, where $k=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{P}(M)-1$, such that
$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p,1)}g_{i}\overline{g_{j}}=\delta_{ij},\quad g_{i}(p)=0\quad\text{ for any }i,j\in\{1,\cdots,k\}.$
and
$\min\limits_{\partial B(p,r)}\sum\limits_{i=1}^{k}|g_{i}|^{2}\geq c(n,\nu)r^{2}$
holds for any $r>0$ and some constant $c=c(n,\nu)$.
Moreover, $(g_{1},\cdots,g_{k})$ serves as an embedding of $M$ onto an affine variety in $\mathbb{C}^{k}$.
Take an $n$-tuple algebraically independent holomorphic functions with minimal degrees in $(g_{1},\cdots,g_{k})$, denoted by $(f_{1},\cdots,f_{n})$. As in section 4 in [16], $(f_{1},\cdots,f_{n})$ serves as a biholomorphism of $M$ onto $\mathbb{C}^{n}$, and any polynomial growth function is a polynomial of them.
Since any $g_{i}$ is a polynomial of these $(f_{1},\cdots,f_{n})$ with degree no larger than $D$ by the three-circle theorem, by a contradiction argument we get
$\min\limits_{\partial B(p,r)}\sum\limits_{i=1}^{n}|f_{i}|^{2}\geq c(n,\nu)r^{\frac{2}{D}}$ (2.2)
holds for any $r>0$ and some constant $c=c(n,\nu)$. In other words, $(f_{1},\cdots,f_{n})$ is proper.
For (3), given any two such global coordinates $(f_{1},\cdots,f_{n})$ and $(h_{1},\cdots,h_{n})$. Assume their degrees are arranged in the nondecreasing order, then the matrix $\left(\frac{\partial h_{i}}{\partial f_{j}}\right)$ and $\left(\frac{\partial f_{i}}{\partial h_{j}}\right)$ are both polynomial matrices of $(f_{1},\cdots,f_{n})$ and are inverse of each other. Therefore they must both be constant matrices. Then due to the normalization condition (2.1), these two matrices are both orthogonal. ∎
In fact, the minimal degrees of all coordinates in Theorem 2.1 (3) described above are just the refined minimal order $\overrightarrow{d_{\operatorname{min}}}$ defined in the introduction, and it has been shown in [16].
Corollary 2.2.
Conjecture 1.5 (1) holds.
Proof.
If $M$ is of nonmaximal volume growth. By Theorem 2 in [12], $\mathcal{O}_{P}(M)=\mathbb{C}$, i.e. $\overrightarrow{d_{\operatorname{min}}}=\overrightarrow{+\infty}$, which is independent of the choice of $p$.
If $M$ is of maximal volume growth, due to [16], there exists a strictly increasing sequence $1\leq d_{1}<d_{2}<\cdots$, so that $\mathcal{O}_{P}(M)=\mathbb{C}\oplus\left(\mathcal{O}_{d_{1}}(M)/\mathbb{C}\right)\oplus\left(\mathcal{O}_{d_{2}}(M)/\mathcal{O}_{d_{1}}(M)\right)\oplus\cdots$ as a complex vector space. For any $k\in\mathbb{N}$, pick a maximal linearly independent vectors $f_{k1},\cdots,f_{km_{k}}$ of $\mathcal{O}_{d_{k}}(M)$ so that they form a basis of $\mathcal{O}_{d_{k}}(M)/\mathcal{O}_{d_{k-1}}(M)$ as quotient of vector spaces and no element in the span is given by polynomials of $\mathcal{O}_{d_{k-1}}(M)$. Then the first $n$ functions $\{f_{11},\cdots,f_{1m_{1}},f_{21},\cdots\}$ form a global coordinate in Theorem 2.1 after normalization as (2.1).
[…]
Remark 2.3.
(1) As can be seen in the proof, if $M$ is of maximal volume growth, Corollary 2.2 still holds without the nonsplitting condition on the universal cover.
(2) For simplicity, if $M$ is of maximal volume growth we just call the coordinate as in (2) in Theorem 2.1 a canonical coordinate on $M$, with degrees $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$. And we arrange them in non-decreasing order of their degrees.
Let $M_{\infty}$ be a tangent cone of infinity of $M$, whose structure has been studied in [14, 15, 9, 19, 18]. Let’s summarize these results in the following:
Theorem 2.4.
Let $(M,g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth. Let $(M_{\infty},p_{\infty},d_{\infty})$ be a tangent cone at infinity of $M$, then:
(1) $(M_{\infty},p_{\infty},d_{\infty})$ is a complex manifold which is biholomorphic to $\mathbb{C}^{n}$. Moreover, it’s a metric Kähler cone with $BK\geq 0$” (see Section 5 in [19] for relevant definitions).
(2) There exists a complete Kähler-Ricci flow solution $h(t)$ on $M_{\infty}$ for $t\in[0,+\infty)$ with nonnegative bisectional curvature such that $(M_{\infty},p_{\infty},d_{\infty})$ is also the Gromov-Hausdorff limit of $h(t)$ as $t\rightarrow 0$. Moreover, $(M_{\infty},h(t))$ is a gradient Kähler-Ricci soliton with nonnegative bisectional curvature.
(3) $\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M_{\infty})=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)$ for any $d\geq 0$. And there exists a n-tuple of polynomial growth functions $(z_{1\infty},\cdots,z_{n\infty})$ serving as a proper biholomorphism onto $\mathbb{C}^{n}$. Each $z_{i\infty}$ is of homogeneous degree $d_{\operatorname{min}}^{(i)}$, i.e. there exists a eigenfunction $\phi_{i}$ of the Laplacian operator on the link such that $z_{i\infty}=r^{d_{\operatorname{min}}^{(i)}}\phi_{i}$ on $M_{\infty}$. They satisfy
$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p_{\infty},1)}z_{i\infty}\overline{z_{j\infty}}=\delta_{ij},\quad z_{i\infty}(p_{\infty})=0\quad\text{ for any }i,j\in\{1,\cdots,n\}.$ (2.3)
Moreover, any element in $\mathcal{O}_{P}(M_{\infty})$ is a polynomial of these functions. In particular, $\mathcal{O}_{P}(M_{\infty})\cong\mathbb{C}[z_{1\infty},\cdots,z_{n\infty}]$.
(4) Any two such global coordinates (consisting of n polynomial growth homothetically homogeneous, holomorphic functions satisfying (2.3), serving as a proper biholomorphism onto $\mathbb{C}^{n}$, and generating $\mathcal{O}_{P}(M_{\infty})$) differ by a constant orthogonal transformation.
(5) Suppose there is a blow-down sequence $(M_{i},g_{i},p_{i})=(M,r_{i}^{-2}g,p)$ converges to $(M_{\infty},d_{\infty},p_{\infty})$ in the Gromov-Hausdorff sense. Let $(z_{1},\cdots,z_{n})$ be a canonical coordinate on $M$. Normalize them to $(z_{1k},\cdots,z_{nk})$ on $M_{k}$ such that
$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p_{k},1)}z_{ik}\overline{z_{jk}}=\delta_{ij},\quad z_{ik}(p_{k})=0\quad\text{ for any }i,j\in\{1,\cdots,n\},k\in\mathbb{N}.$ (2.4)
Then $(z_{1k},\cdots,z_{1n})$ sub-sequentially converges to a global coordinate $(z_{1\infty},\cdots,z_{n\infty})$ as in (3) on $M_{\infty}$ uniformly on compact subsets as $k\rightarrow+\infty$.
(6) The volume of the unit ball in $M_{\infty}$ is $\frac{\omega_{2n}}{d_{\operatorname{min}}^{(1)}\cdots d_{\operatorname{min}}^{(n)}}$. In particular, Conjecture 1.5 (2) holds.
Proof.
(1) and (2) were proved in Proposition 6.1 in [19].
For (3), for any $d\geq 0$, $\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M_{\infty})=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)$ by Proposition 4.1 in [16]. Now we show the existence of such coordinate.
Firstly we claim that the homothetic vector field $r\frac{\partial}{\partial r}$ is holomorphic on $M_{\infty}$. By Theorem 1.9 there exists a long-time Kähler-Ricci flow solution $g(t)$ on $M$ such that $(M_{\infty},J_{\infty},p_{\infty},h(t))$ is the pointed Cheeger-Hamilton limit of a blow-down sequence of $g(t)$, i.e. $(M,J,p,g_{i}(t)=\frac{1}{t_{i}}g(t_{i}t)$ for some sequence $t_{i}\rightarrow+\infty$. In other words, $J$ converges to $J_{\infty}$ through a sequence of diffeomorphisms exhausted on $M_{\infty}$. By Claim 4.3 in [16], the complexification of $r\frac{\partial}{\partial r}$ is in the span of several holomorphic vector fields on $M_{\infty}$. Moreover, $(r\frac{\partial}{\partial r})\mathcal{O}_{d}(M_{\infty})\subset\mathcal{O}_{d}(M_{\infty})$ holds for any $d\geq 0$. The claim is confirmed.
Therefore, $r\frac{\partial}{\partial r}$ is a contracted holomorphic field on $M_{\infty}$, from a result in [25], $M_{\infty}$ is biholomorphic to $\mathbb{C}^{n}$. Now follow Section 4 in [16], replace the holomorphic vector field $X$ in the proof by $r\frac{\partial}{\partial r}$, we conclude the proof of (3) except the homogenity of these coordinates. This is also obvious since we can just choose the highest order terms of each coordinate functions and they certainly form a new biholomorphism onto $\mathbb{C}^{n}$.
The proof of (4) is the same as the proof of (3) in Theorem 2.1.
For (5), the mean value inequality [10] implies that,
$M_{z_{ik}}(\frac{1}{2})=\sup\limits_{B(p_{k},\frac{1}{2})}|z_{ik}|\leq C(n),$
holds for any $1\leq i\leq n$ and $k\in\mathbb{N}$. By the three-circle theorem in [13] and Cheng-Yau gradient estimate [5],
$M_{z_{ik}}(r)\leq C(n)r^{d_{\operatorname{min}}^{(i)}},\quad|dz_{ik}|\leq C(n)r^{d_{\operatorname{min}}^{(i)}-1}$
on $B_{p_{i}}(r)$ for any $1\leq i\leq n$, $k\in\mathbb{N}$ and $r\geq\frac{1}{2}$. Arzela-Ascoli theorem implies that $z_{ik}$ converges to a holomorphic function $z_{i\infty}$ of homogeneous degree no more than $d_{\operatorname{min}}^{(i)}$ uniformly on compact subsets. By (2.2) and the Gromov-Hausdorff convergence,
$\min\limits_{\partial B(p_{\infty},r)}\sum\limits_{i=1}^{n}|z_{i\infty}|^{2}\geq c(n,\nu)r^{\frac{2}{D(n,\nu)}}>0.$ (2.5)
Firstly, we show these functions are algebraically independent. Suppose not, then there exists a nontrivial polynomial $P\in\mathbb{C}[x_{1},\cdots,x_{n}]$ such that
$P(z_{1\infty},\cdots,z_{n\infty})\equiv 0$
on $M_{\infty}$. Lift to $M_{k}$ for large $k$ and use the properness (2.4) of $(z_{1k},\cdots,z_{nk})$, it implies $|P(z_{1k},\cdots,z_{nk})|$ can be arbitrarily small on arbitrarily large ball in $\mathbb{C}^{n}$. Thus $P\equiv 0$. And $z_{i\infty}$ must of homogeneous degree $d_{\text{min}}^{(i)}$ by the dimension estimate in (3), because on $M_{\infty}$ the minimal degree of a n-tuple algebraically independent holomorphic homogeneous functions has to be $\overrightarrow{d_{\operatorname{min}}}$. Therefore, (4) implies that $(z_{1\infty},\cdots,z_{n\infty})$ must form a global coordinate as in (3).
For (6), the first statement was proved in the proof of Theorem 3 in page 13 in [18]. This concludes the proof of Conjecture 1.5 (2) in the maximal volume growth case. The case of nonmaximal volume growth is trivial. ∎
Remark 2.5.
(1) As can be seen in the proof, if $M$ is of maximal volume growth, Conjecture 1.5 (2) still holds without the nonsplitting condition on the universal cover.
(2) For simplicity, we just call the coordinate as in (2) in Theorem 2.4 a canonical coordinate on $M_{\infty}$, with degrees $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$. And we arrange them in non-decreasing order of their degrees.
A natural corollary is the following homogeneity property of $\mathcal{O}_{P}(M)$:
Corollary 2.6.
Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth. Then for any $f,g\in\mathcal{O}_{P}(M)$,
$\operatorname{deg}(fg)=\operatorname{deg}(f)+\operatorname{deg}(g).$
Proof.
Take a canonical coordinate $(f_{1},\cdots,f_{n})$ on $M$ with
$f_{i}(p)=0\quad\text{ for any }i\in\{1,\cdots,n\}.$
For any sequence $r_{k}\rightarrow+\infty$, on $(M_{k},p_{k})=(r^{-2}_{k}M,p)$, we can normalize them such that
$f_{ik}=\frac{f_{i}}{M_{ik}},\quad\text{ where }M_{ik}=\sup\limits_{B_{g}(p,r_{k})}|f_{i}|$
for any $i\in\{1,\cdots,n\}$ and $k\in\mathbb{N}$. Therefore on each $M_{k}$,
$\sup\limits_{B(p_{k},1)}|f_{ik}|=1.$
Suppose $(M_{i},p_{i})\rightarrow(M_{\infty},p_{\infty})$ in the sense of Gromov-Hausdorff convergence, as in Proposition 5 in [15], each $f_{ik}\rightarrow f_{i\infty}$ with $f_{i\infty}$ is of homoegeneous degree $d_{\operatorname{min}}^{(i)}$.
As these functions generate $\mathcal{O}_{P}(M)$, we just need to prove: for any $f=f_{1}^{k_{1}}\cdots f_{n}^{k_{n}}$, $\operatorname{deg}(f)=\sum\limits_{i=1}^{n}k_{i}d_{\operatorname{min}}^{(i)}$.
Now let $f=f_{1}^{k_{1}}\cdots f_{n}^{k_{n}}$, by the three-circle theorem, for any $a>1$
$\lim_{r\to\infty}\frac{M_{f}(ar)}{M_{f}(r)}=a^{\operatorname{deg}(f)}.$
On the other hand,
$\lim_{r\to\infty}\frac{M_{f}(ar)}{M_{f}(r)}$ $=\lim\limits_{r\to\infty}\frac{M_{f_{1}^{k_{1}}\cdots f_{n}^{k_{n}}}(ar)}{M_{f_{1}^{k_{1}}\cdots f_{n}^{k_{n}}}(r)}$
$=\lim\limits_{k\to\infty}\frac{M_{f_{1}^{k_{1}}\cdots f_{n}^{k_{n}}}(ar_{k})}{M_{f_{1}^{k_{1}}\cdots f_{n}^{k_{n}}}(r_{k})}$
$=\lim\limits_{k\to\infty}\frac{M_{f_{1k}^{k_{1}}\cdots f_{nk}^{k_{n}}}(a)}{M_{f_{1k}^{k_{1}}\cdots f_{nk}^{k_{n}}}(1)}$
$=\frac{M_{f_{1\infty}^{k_{1}}\cdots f_{n\infty}^{k_{n}}}(a)}{M_{f_{1\infty}^{k_{1}}\cdots f_{n\infty}^{k_{n}}}(1)}$
$=a^{\sum\limits_{i=1}^{n}k_{i}d_{\operatorname{min}}^{(i)}}$
Compare these we complete the proof. ∎
Recently, Chu-Hao [7] obtain an optimal rigidity of the dimension estimate for polynomial growth holomorphic functions. Now we give an alternative proof, which seems much simpler.
Theorem 2.7.
Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature. Suppose there exists $d\geq 1$ such that
$\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)>\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(\mathbb{C}^{n})-\binom{n+d-2}{d-1}$,
then $M$ is biholomorphically isometric to $\mathbb{C}^{n}$ with Euclidean metric.
Proof.
Suppose $M$ is of maximal volume growth, then take $(f_{1},\cdots,f_{n})$ be a canonical coordinate with minimal degree $(d^{(1)}_{\operatorname{min}},\cdots,d^{(n)}_{\operatorname{min}})$ with $d^{(i)}_{\operatorname{min}}\geq 1$ in non-decreasing order.
From Corollary 2.6, $\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)=\#\{(m_{1},\cdots,m_{n})\in\mathbb{Z}^{n}_{\geq 0}\mid\sum\limits_{i=1}^{n}m_{i}d^{(i)}_{\min}\leq d\}$. We know that $M$ is biholomorphically isometric to $\mathbb{C}^{n}$ if and only if $d^{(i)}_{\min}=1$ for all $i$. Suppose $M$ is not isometric to $\mathbb{C}^{n}$. Then $d^{(n)}_{\min}>1$, then
$\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)$ $=\#\{(m_{1},\cdots,m_{n})\in\mathbb{Z}^{n}_{\geq 0}\mid\sum\limits_{i=1}^{n}m_{i}d^{(i)}_{\min}\leq d\}$
$\leq\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(\mathbb{C}^{n})-\#\{(m_{1},\cdots,m_{n-1})\in\mathbb{Z}^{n-1}_{\geq 0}\mid\sum\limits_{i=1}^{n}m_{i}<d\}$
$=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(\mathbb{C}^{n})-\binom{n+d-2}{d-1}$.
with equality holds if and only $d^{(i)}_{\operatorname{min}}=1$ for $i\leq n-1$ and $1<d^{(n)}_{\operatorname{min}}<\frac{d}{d-1}$. This contradicts with our assumption.
Now suppose the universal cover $\widetilde{M}$ splits. By the proof of Corollary 7.1 in [11], $\widetilde{M}=M_{1}^{m_{1}}\times M_{2}^{n-m_{1}}$, where $M_{1},M_{2}$ are both simply connected and has nonnegative bisectional curvature. Also $M_{1}$ is of maximal volume growth. $\mathcal{O}_{d}(M)$ can be seen as the $G$-invariant subring of $\mathcal{O}_{d}(M_{1})$, $G$ is a compact Lie group acting on $M_{1}$.
Therefore
$\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(\mathbb{C}^{n})-\binom{n+d-2}{d-1}<\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)\leq\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M_{1}),$
From the paragraph above, $M_{1}$ must be biholomorphically isometric to $\mathbb{C}^{n}$, i.e. $\widetilde{M}=\mathbb{C}^{n}$. And the $G$-action on $\mathbb{C}^{n}$ must be the deck transformation of $\pi_{1}(M)$.
Claim: Let $\pi:\widetilde{M}\rightarrow M$ be the covering map, $f\in\mathcal{O}_{P}(M)$, then $\operatorname{deg}(f)=\operatorname{deg}(f\circ\pi)$.
It’s obvious since $\pi(B(\tilde{p},r))=B(p,r)$ for any $\pi(\tilde{p})=p$ and any $r>0$.
Since $\mathcal{O}_{P}(M)$ is finitely generated, take a family of minimal generators of it, say $\{f_{1},\cdots,f_{N}\}\subset\mathcal{O}_{P}(\mathbb{C}^{n})$. By the claim we know that $\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)=\#\{\text{monomial of }f_{1},\cdots,f_{N}\text{ which is in }\mathcal{O}_{d}(\mathbb{C}^{n})\}$.
Similar argument as above shows that
$\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)\leq\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(\mathbb{C}^{n})-\binom{n+d-2}{d-1}$
unless $\{f_{1},\cdots,f_{N}\}=A\{z_{1},\cdots,z_{n}\}+b$ for some constant $A,b$, in which case $\pi_{1}(M)$ is trivial. Therefore we complete the proof. ∎
Remark 2.8.
Similar argument can also prove the splitting theorem in Section 3 in [7], by using Theorem 4.1 in [23].
2.2. Gradient expanding Kähler-Ricci solitons
Definition 2.9.
A Kähler-Ricci soliton consists of a triple $(M,g,X)$, where $M$ is a Kähler manifold, $X$ is a holomorphic vector field on $M$, and $g$ is a complete Kähler metric on $M$ whose Kähler form $\omega$ satisfies
$\operatorname{Ric}-\frac{1}{2}\mathcal{L}_{X}\omega+\lambda\omega=0$
for some $\lambda\in\{-1,0,1\}$. A Kähler-Ricci soliton is said to be expanding if $\lambda=1$. Here $X$ is called the soliton vector field. If in addition, $X=\nabla^{g}f$ for some real-valued smooth function $f$ on $M$, then we say $(M,g,X)$ is a gradient Kähler-Ricci soliton. In this case, we call $f$ the potential function of the soliton. It’s equivalent to say
$\operatorname{Ric}_{i\bar{j}}-f_{i\bar{j}}+\lambda g_{i\bar{j}}=0,\quad f_{ij}=f_{\bar{i}\bar{j}}=0.$
They we list some important properties of gradient expanding Kähler-Ricci solitons with nonnegative Ricci curvature which can be found in [2, 6, 30, 4, 28].
Proposition 2.10.
Let $(N^{n},J,g,f)$ be a complete noncompact gradient expanding Kähler-Ricci solitons with nonnegative Ricci curvature. Then
(1) $R+|\nabla^{g}f|^{2}-f$ is constant on $N$, therefore we always normalize $f$ so that $f=R+|\nabla^{g}f|^{2}$.
(2) The potential $f$ is a strictly convex exhaustion function with $\nabla^{g}f$ has the unique zero at $O\in N$, and $R$ attains its maximum at $O$. In particular, if $N$ has nonnegative bisectional curvature, then $N$ has bounded curvature.
(3) $(N,g)$ is biholomorphic to $\mathbb{C}^{n}$ and has maximal volume growth.
(4) $\nabla^{g}f$ is a complete vector field. Let $\varphi(t)$ be a family of biholomorphism of $(N,J)$ with
$\frac{\partial}{\partial t}\varphi(t-1)(x)=-\frac{1}{t}\nabla^{g}f(\varphi(t-1)(x)),\varphi(0)=\text{id}.$ (2.6)
It follows that
$g(t)=t\varphi(t-1)^{*}g$ (2.7)
solves the Kähler-Ricci flow equation on $[0,+\infty)$ with $g(1)=g$.
(5) ((13) in [28]) The potential function $f$ satisfies the estimate
$f(\varphi(O,t))\leq f(O)e^{2t}.$ (2.8)
3. Proof of Theorem B
The key observation is the following which may be well-known for experts in the fields. However, for reader’s convenience, we give a proof here.
Lemma 3.1.
Let $(N,J,h,O,f)$ be a complete noncompact expanding Kähler-Ricci soliton with nonnegative bisectional curvature and we normalize it by assuming $f=R+|\nabla^{h(1)}f|^{2}$. Let
$h(t)=t\varphi(t-1)^{*}h$
be the unique Kähler-Ricci flow solution on $[0,+\infty)$ such that $h(1)=h$. Where
$\frac{\partial}{\partial t}\varphi(t-1)(x)=-\frac{1}{t}\nabla^{g}f(\varphi(t-1)(x)),\varphi(0)=\text{id}.$
Then $(N,h(1))$ has a unique tangent cone at infinity $(N_{\infty},h_{\infty},O_{\infty})$. Moreover, $(N,h(t),O)\rightarrow(N_{\infty},h_{\infty},O_{\infty})$ as $t\rightarrow 0$ in the Gromov-Hausdorff sense.
Proof.
Replace $(M,g)$ in Theorem 1.6 by $(N,h(1))$, we conclude that there exists a metric Kähler cone $N_{\infty}$ such that $(N,h(t),O)\rightarrow(N_{\infty},h_{\infty},O_{\infty})$ as $t\rightarrow 0$ in the Gromov-Hausdorff sense. So it remains to prove this cone $N_{\infty}$ is the unique tangent cone of $(N,h(1))$. In fact we have
$(N_{\infty},h_{\infty},O_{\infty})$ $=\lim\limits_{t\to 0}(N,h(t),O)$
$=\lim\limits_{t\to 0}(N,t\varphi(t-1)^{\star}h(1),O)$
$=\lim\limits_{t\to 0}(N,th(1),\varphi(t-1)(O))$
$=\lim\limits_{t\to 0}(N,th(1),(O)).$
All the limits above are in the sense of Gromov-Hausdorff convergence. The second-to-last equality holds by the following claim:
Claim 1.
$\lim\limits_{t\to 0}d_{th(1)}(\varphi(t-1)(O),O)=0$.
Proof.
For any $s\leq 1$, by (2) and (2.8) in Proposition 2.10,
$|\nabla^{h(1)}f|^{2}\left(\varphi(s)(O)\right)\leq R+|f(\varphi(s)(O))|\leq C(1+e^{2})\leq C.$
Integrating along $\varphi(s)(O)$, we obtain
$\sqrt{t}d_{h(1)}(\varphi(t-1)(O),O)\leq\sqrt{t}\int_{t}^{1}\frac{C}{s}ds=-C\sqrt{t}\log(t)\rightarrow 0\text{ as }t\rightarrow 0.$
Thus, we now have a clear geometric framework characterizing the long-time asymptotics of the Kähler-Ricci flow on a complete noncompact Kähler manifold with bounded nonnegative bisectional curvature and maximal volume growth, along with its relationship to its tangent cones.
Now we prove Theorem B.
Proof of Theorem B:
Following the notation of Theorem 1.6, take the Poincaré coordinate $(w_{1},\cdots,w_{n})$ as in page 2622 in [28] on $(N,h(1),O)$ and normalize them such that
$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B_{h(1)}(O,1)}{w_{i}}\overline{w_{j}}=\delta_{ij},\quad w_{i}(O)=0\quad\text{ for any }i,j\in\{1,\cdots,n\}.$
According to the proof of Theorem 1.4 in [28], the aforementioned Poincaré coordinate $(w_{1},\cdots,w_{n})$ is a canonical coordinate on $(N,h(1),O)$. Following the procedure of (5) in Theorem 2.4 on $(N,h(1),O)$ and $(M,g,p)$, we get two canonical coordinates $(z_{1\infty},\cdots,z_{n\infty})$ and $(w_{1\infty},\cdots,w_{n\infty})$ on $N_{\infty}=M_{\infty}$, with degree $\overrightarrow{d_{\operatorname{min}}}(M)$ and $\overrightarrow{d_{\operatorname{min}}}(N)$. Thus from (4) in Theorem 2.4,
$\overrightarrow{d_{\operatorname{min}}}(M)=\overrightarrow{d_{\operatorname{min}}}(N)=\{\mu_{1}+1,\cdots,\mu_{n}+1\}.$
Therefore, (1),(2) has been proved. For (3), note that in [18], Liu has proved that $\operatorname{ASCD}(M,g(t))=4n(\sum\limits_{i=1}^{n}d^{(i)}_{\operatorname{min}}-n)=4n(\sum\limits_{i=1}^{n}\mu_{i})$ is invariant along $t$. The proof is complete.
As an application, we can use it to prove Corollary 1.11:
Proof of Corollary 1.11:
Firstly, we assume $(M,g)$ has bounded curvature. Take a canonical coordinate $(f_{1},\cdots,f_{n})$ on $(M,g)$, with degree $\overrightarrow{d_{\operatorname{min}}}$. By comparing the distance function $d_{g(t)}(x,p)$ along the flow using the shrinking balls lemma by Simon and Topping (Corollary 3.3 in [26]), we may conclude that $(f_{1},\cdots,f_{n})$ is still a global coordinate on $(M,g(t))$, with the same degree $\overrightarrow{d_{\operatorname{min}}}$. Let $g_{k}(t)=\frac{1}{t_{i}}g(t_{i}t)$, we normalized them such that
$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B_{g_{k}(1)}(O,1)}{f_{ik}}\overline{f_{jk}}=\delta_{ij},\quad f_{ik}(p)=0\quad\text{ for any }i,j\in\{1,\cdots,n\},k\in\mathbb{N}.$
which makes $(f_{1k},\cdots,f_{nk})$ serving as a canonical coordinate on $(M,g_{i}(1))$.
Following the same proof as in (5) in Theorem 2.4, these $(f_{1k},\cdots,f_{nk})$ converges smoothly to $(f_{1\infty},\cdots,f_{n\infty})$ on $(N,h(1))$ and serving as a canonical coordinate on $N$. Due to (3) in Theorem 2.1, WLOG we can assume $(f_{1\infty},\cdots,f_{n\infty})$ is just a Poincaré coordinate on $(N,h(1))$. Therefore
$\operatorname{Ric}_{h(1)}\left(\frac{\frac{\partial}{\partial f_{i\infty}}}{|\frac{\partial}{\partial f_{i\infty}}|},\overline{\frac{\frac{\partial}{\partial f_{i\infty}}}{|\frac{\partial}{\partial f_{i\infty}}|}}\right)=\mu_{i}.$
Finally, by the smooth convergence of $g_{i}(1)$ and the proof of Theorem 1.2 in [3], the Corollary is proved. In fact, for any $i$ and any nonzero $v\in V_{1}\oplus\cdots\oplus V_{i}$ but $v\notin V_{1}\oplus\cdots\oplus V_{i-1}$,
$\lim\limits_{t\to\infty}t\operatorname{Ric}(v(t),\overline{v(t)})=\mu_{i}$
always holds.
If $(M,g)$ has unbounded curvature, thanks to Theorem 1.9, above argument can still be applied.
4. Proof of Theorem A
Proof of Theorem A:
Firstly we assume $M$ is of nonmaximal volume growth, assume there exists a nonzero polynomial holomorphic $n$-form $s\in H^{0}_{P}(M,\mathcal{K}_{M})$, the Poincaré-Lelong equation states that
$\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\log\|s\|^{2}=(s)+\operatorname{Ric}\geq 0,$
where $(s)$ is the zero divisor of $s$. In other words, $h=\log(\|s\|^{2}+1)$ is a smooth plurisubharmonic function of logarithmic growth. Since the universal cover of $M$ does not split, the heat-flow method of Ni-Tam [23] can be applied to perturb $h$ to $h^{\prime}$ which is a smooth plurisubharmonic function of still logarithmic growth (for details, see Theorem 0.1 (iii) in [23]). Then we solve the $\bar{\partial}$-equation by applying the Hörmander $L^{2}$-estimate similarly as the proof of Lemma 3 in [12], there exists a nonconstant polynomial growth holomorphic function on $M$, contradicting with Theorem 2 in [12]. As a result, $D_{\operatorname{min}}=+\infty=\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n$.
Now we assume $M$ is of maximal volume growth. Let $(z_{1},\cdots,z_{n})$ be a canonical coordinate of $M$ with degrees $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$ as stated in Remark 2.3. By gradient estimate [5],
$\operatorname{deg}(dz_{1}\wedge\cdots\wedge dz_{n})\leq\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n,$
therefore $D_{\operatorname{min}}\leq\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n$.
The proof of the opposite direction is similar to the proof of Corollary 1.5 in [28], but a little more complicated. Let $s\in H^{0}_{D_{\operatorname{min}}}(M,\mathcal{K}_{M})$, then note that Corollary 3 in [13] also holds for holomorphic sections of holomorphic line bundle with nonpoaitive curvature, there exists a constant $C>0$ such that
$M_{s}(r,0)\leq Cr^{D_{\operatorname{min}}}\quad\text{ for any }r>0.$ (4.1)
Fix $t_{k}\rightarrow+\infty$, there exists a long-time Kähler-Ricci flow solution $g(t)$ which satisfies properties in Theorem 1.9. And set $g_{k}(t)=\frac{1}{t_{i}}g(t_{i}t)$, assume $t_{k}\geq 1$.
Claim 2.
[…]
Proof.
By the shrinking balls lemma by Simon and Topping (Corollary 3.3 in [26]), for any $t_{k}>0$, we have
$d_{g}\leq d_{g(t_{k})}+C^{\prime}_{k}\sqrt{t_{k}}$
on $M\times[0,+\infty)$.
Also, thank to Corollary 1.10, (1.2) also holds for $M$ with unbounded curvature, which means that
$\lim\limits_{t\to\infty}\log\frac{\det(g_{i\bar{j}}(x,t))}{\det(g_{i\bar{j}}(x,0))}=-C.$
for some nonnegative constant $C$ which relies only on $M$ and is independent of $x$. Also by the Ricci flow equation,
$\frac{d}{dt}\log\frac{\det(g_{i\bar{j}}(x,t))}{\det(g_{i\bar{j}}(x,0))}=-R(x,t)\leq 0.$
So
$\log\frac{\det(g_{i\bar{j}}(x,t))}{\det(g_{i\bar{j}}(x,0))}\geq-C$
holds uniformly on $M\times[0,+\infty)$.
Therefore, on $(M,g(t_{k}))$, we have
$M_{s}(r,t_{k})\leq M_{s}(r+C^{\prime}_{k}\sqrt{t_{k}},0)\frac{\det(g_{i\bar{j}}(x,0))}{\det(g_{i\bar{j}}(x,t))}\leq C_{k}r^{D_{\operatorname{min}}}.$
The claim is proved. ∎
[…]
Thanks to Theorem 1.9, the boundedness condition of curvature in Theorem 1.6 can be removed. Using the notations in Theorem 1.6, recall the sequence $(U_{k},F_{k}^{*}J.F_{k}^{*}(g_{k}(1)),O)$ converges smoothly to $(N,J_{\infty},h(1),O)$ uniformly on compact subsets. Define
$s_{k}=\frac{s\mid_{F_{k}(U_{k})}}{\sup\limits_{B_{g_{k}(1)}(p,1)}\|s\|},\quad\tilde{s}_{k}=F_{k}^{*}s_{k}.$ (4.2)
Since $\mathcal{K}_{M}(g_{k}(1))$ has nonpositive curvature, three-circle theorem in [13] can be applied to get for any fixed $r>0$, for large $k$,
$\sup\limits_{B_{g_{k}(1)}(p,r)}\|s_{k}\|\leq r^{D_{\operatorname{min}}}.$
After pulling back by $F_{k}$, we obtain
$\sup\limits_{B_{F_{k}^{*}g_{k}(1)}(O,r)}\|s_{k}\|\leq r^{D_{\operatorname{min}}}.$ (4.3)
holds for large $k$.
Fix a nonvanishing holomorphic $n$-form $\Omega$ on $N$, then write $s_{k}=f_{k}\Omega_{k}$ on $U_{k}$, where $f_{k}$ is a holomorphic function with respect to $F_{k}^{*}J$ on $U_{k}$, $\Omega_{k}$ is a nonvanishing holomorphic $n$-form with respect to $F_{k}^{*}J$ and converges to $\Omega$ uniformly on each fixed compact subset in $N$ (For example, take $U_{k}$ be a sequence of exhausted Euclidean balls in $N\cong\mathbb{C}^{n}$).
Therefore there exists a constant $C=C(K)$ such that $\|\Omega_{k}\|_{F_{k}^{*}(g_{k}(1))}\geq C(K)$ uniformly on $K\subset N$ compact. Then we can deduces
$\sup\limits_{K}|f_{k}|\leq\frac{1}{C(K)}r^{D_{\operatorname{min}}}$
holds for all large $k$. Gradient estimate and Arzela-Ascoli theorem imply that $f_{k}\rightarrow f$ uniformly on compact subsets in $N$, and $f$ is holomorphic with respect to $J_{\infty}$. In other words, sub-sequentially $s_{k}$ converges to a holomorphic $n$-form $s^{\prime}=f\Omega$ on $N$.
From (4.2) and (4.3), $s^{\prime}$ is nonzero ($\sup\limits_{B_{h(1)}(O,1)}\|s^{\prime}\|=1$) and satisfies
$\sup\limits_{B_{h(1)}(O,r)}\|s_{k}\|\leq r^{D_{\operatorname{min}}}.$
Combining with (1.5) we obtain
$D_{\operatorname{min}}\geq D_{\operatorname{min}}(N)=\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}(N)-n=\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n.$
Therefore
$D_{\operatorname{min}}=\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n.$
A much simpler way is consider directly a tangent cone of $M$. Assume $(M,g_{i}=r_{i}^{-2}g)\rightarrow M_{\infty}$ in the Gromov-Hausdorff sense, where $r_{i}\rightarrow\infty$. Take a canonical coordinates $(z_{1},\cdots,z_{n})$ on $(M,g)$, we know that
$D_{\operatorname{min}}=\operatorname{deg}(dz_{1}\wedge\cdots\wedge dz_{n}).$
And from (4) we have known that
$D_{\operatorname{min}}\leq\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n.$
By the three-circle theorem, assume
$D_{\operatorname{min}}\leq\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n-\varepsilon$
for some $\varepsilon>0$.
Following the notations in Theorem 2.3 (5), and define
$\varphi_{i}=\log\|dz_{1i}\wedge\cdots\wedge dz_{ni}\|^{2}_{g_{i}}$
As in [18], since $\varphi_{i}$ are psh on $M_{i}$, by passing to a subsequence, we may assume $\varphi_{i}$ converges to a psh function on $M_{\infty}$ in $L_{\operatorname{loc}}^{1}$, denoted by $\varphi=\log|dz|^{2}$. Therefore $e^{\varphi_{i}}\rightarrow e^{\varphi}$ almost everywhere.
By the three circle theorem as in the proof of Theorem 2.3 (5), there exists a $C$ such that
$e^{\varphi_{i}}\leq Cr_{i}^{2D_{\operatorname{min}}}\leq Cr_{i}^{2(\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n-\varepsilon)}.$
uniformly. Thus $e^{\varphi}\leq Cr^{2(\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n-\varepsilon)}$ almost everywhere, which contradicts with Claim 1 in [18] if $\varepsilon>0$, since $e^{\varphi}$ is homogeneous of degree $2(\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n)$. This directly implies that $D_{\operatorname{min}}=\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n.$

### v2

@@
 Our main results are: 1.
-Precise estimates among refined minimal degree of polynomial growth holomorphic functions and holomorphic volume forms, $\operatorname{AVR}$ (asymptotic volume ratio) and $\operatorname{ASCD}$ (average of scalar curvature decay) are established.
+Precise relations among refined minimal degree of polynomial growth holomorphic functions and holomorphic volume forms, $\operatorname{AVR}$ (asymptotic volume ratio) and $\operatorname{ASCD}$ (average of scalar curvature decay) are established.
 2.
 The Lyapunov asymptotic behavior of the Kähler-Ricci flow can be described in terms of polynomial growth holomorphic functions.
-This provides a unifying perspective that bridges the two distinct proofs of Yau’s uniformization conjecture.
+This provides a unifying perspective that bridges the two distinct proofs of Yau’s uniformization conjecture by Liu and Chau-Lee-Tam.
 These resolve two conjectures made by Yang.
+3.
+the link of any tangent cone is homeomorphic to a sphere, providing a Kähler analogue of a result of Kapovitch in the Riemannian case.
 1. Introduction
 As part of Yau’s program to study complex manifolds of parabolic type, he proposed the following well-known uniformization conjecture in 1970s:
-Conjecture 1.1 (Yau’s Uniformization conjecture, [29]).
+Conjecture 1 (Yau’s Uniformization Conjecture, [Yau91]).
 A complete noncompact Kähler manifold $(M^{n},g)$ with positive bisectional curvature is biholomorphic to $\mathbb{C}^{n}$.
-In the maximal volume growth case, Liu’s breakthrough [16] confirmed the conjecture by combining Gromov-Hausdorff convergence techniques with the three-circle theorem developed in [13].
-An alternative proof, based on the Kähler-Ricci flow, was later provided by Lee-Tam [8], building on results in Chau-Tam [3].
-These advances led to a deeper understanding of the structure of such manifolds.
-Notably, through the works of [20, 21, 24, 11, 12, 17], the following conjecture of Ni has been established as a theorem:
-Theorem 1.2 (Corollary 3.2 in [20], Theorem 1.2 in [24], Theorem 2 in [12], Theorem 1.4 in [11], Corollary 2.16 in [17]).
+In the maximal volume growth case, Liu’s breakthrough [Liu19] confirmed the conjecture by combining Gromov-Hausdorff convergence techniques with the three-circle theorem developed in [Liu16c].
+An alternative proof, based on the Kähler-Ricci flow, was later provided by Lee-Tam [LT20], building on results in Chau-Tam [CT06].
+These advances lead to a deeper understanding of the structure of such manifolds.
+Notably, through the works of [Ni04, Ni05, NT13, Liu16a, Liu16b, Liu21], the following conjecture of Ni has been established as a theorem:
+Theorem 1 (Corollary 3.2 in [Ni04], Theorem 1.2 in [NT13], Theorem 2 in [Liu16b], Theorem 1.4 in [Liu16a], Corollary 2.16 in [Liu21]).
 Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature. Assume that the universal cover of $M$ does not split. Then the following conditions are equivalent:
 (1) $M$ is of maximal volume growth, i.e.
-$\operatorname{AVR}(M,g)=\lim_{r\to\infty}\frac{\operatorname{Vol}\left(B(p,r)\right)}{\omega_{2n}r^{2n}}=\nu>0.$
+$\operatorname{AVR}(M,g)=\lim_{r\to\infty}\frac{\operatorname{Vol}\left(B(p,r)\right)}{\omega_{2n}r^{2n}}>0.$
 Here $\omega_{2n}$ is the volume of the unit ball in $\mathbb{C}^{n}$.
@@
 (3) The average scalar curvature decay is finite, i.e.
-$\operatorname{ASCD}(M,g)=\limsup\limits_{r\to\infty}r^{2}\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p,r)}S$
+$\operatorname{ASCD}(M,g)=\limsup\limits_{r\to\infty}r^{2}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p,r)}S$
 is finite. Here $S$ is the scalar curvature.
-Remark 1.3.
+Remark.
 (1) Note that both $\operatorname{AVR}(M,g)$ and $\operatorname{ASCD}(M,g)$ are independent of the choice of $p$.
-(2) A very recent result of Liu [18] implies that for any $(M,g)$ with nonnegative bisectional curvature, the $\limsup$” in the definition of $\operatorname{ASCD}(M,g)$ above can be replaced by $\lim$”, see also [22].
-In his thesis [27], Yang defined:
-Definition 1.4.
+(2) A very recent result of Liu [Liu24] shows that for any $(M,g)$ with nonnegative bisectional curvature, the $\limsup$” in the definition of $\operatorname{ASCD}(M,g)$ above can be replaced by $\lim$”, see also [Ni12].
+In his thesis [Yang13], Yang defined:
+Definition 1.
 Let $(M^{n},g)$ be a complete Kähler manifold with nonnegative bisectional curvature, for a fixed $p\in M$, define
@@
 […]
-where $\mathcal{K}_{M}$ is the canonical line bundle on $M$, $\mathcal{O}_{P}(M)$ is the space of all polynomial growth holomorphic functions on $M$ and $H^{0}_{P}(M,\mathcal{K}_{M})$ is the space of all polynomial growth holomprhic $n$-forms on $M$.
-Set $d_{\operatorname{min}}=+\infty$ or $D_{\operatorname{min}}=+\infty$ if $M$ does not admit any nonconstant holomorphic functions of polynomial growth or $\mathcal{K}_{M}$ admits no nonzero holomorphic section of polynomial growth.
-Now assume $(M^{n},\omega)$ satisfies the assumption in Theorem 1.2 and of maximal volume growth.
-Since the Kodaira dimension $K(M^{n})=n$, $\mathcal{O}_{P}(M)$ is “holomorphically regular” in the sense that we can always find “local coordinate by global polynomial growth functions”.
-The refined minimal degree can be also defined in [27] by
+where $\mathcal{K}_{M}$ is the canonical line bundle on $M$, $\mathcal{O}_{P}(M)$ is the space of all polynomial growth holomorphic functions on $M$ and $P(M,\mathcal{K}_{M})$ is the space of all polynomial growth holomorphic $n$-forms on $M$.
+Set $d_{\operatorname{min}}=+\infty$ or $D_{\operatorname{min}}=+\infty$ if $M$ does not admit any nonconstant holomorphic functions of polynomial growth or $\mathcal{K}_{M}$ admits no nonzero holomorphic sections of polynomial growth.
+Now assume $(M^{n},\omega)$ satisfies the assumption in Theorem 1 and is of maximal volume growth.
+Since the Kodaira dimension $K(M)=n$, $\mathcal{O}_{P}(M)$ is “holomorphically regular” in the sense that we can always find “local coordinate by global polynomial growth holomorphic functions”.
+The refined minimal degree can be also defined in [Yang13] by
 $\overrightarrow{d_{\operatorname{min}}}(p):=\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right):=\inf_{\{f_{1},\cdots,f_{n}\}}\left\{\limsup_{x\rightarrow\infty}\frac{\log|f_{1}(x)|}{\log d(x,p)},\cdots,\limsup_{x\rightarrow\infty}\frac{\log|f_{n}(x)|}{\log d(x,p)}\right\}$
-where the infimum is taken among any $n$-tuple of global holomorphic functions that gives local coordinates at $p$ with the corresponding $\limsup_{x\rightarrow\infty}\frac{\log|f_{i}(x)|}{\log d(x,p)}$ arranged in a non-decreasing order for $1\leq i\leq n$, we denote this set by $\mathcal{O}_{P,p}(M)$.
-In other words, for any $1\leq k\leq n$ we have $d_{\operatorname{min}}^{(k)}=\inf_{f_{k}}\limsup_{x\to\infty}\frac{\log|f_{k}(x)|}{\log d(x,p)}$, where the infimum is taken among all possible $f_{k}$ that appears in the $k$-th component of some sequence in $\mathcal{O}_{P,p}(M)$.
-Note that apriori it’s unclear if $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$ can be obtained by an n-tuple holomorphic functions in $\mathcal{O}_{P,p}(M)$.
+where the infimum is taken among any $n$-tuple of global holomorphic functions that gives local coordinate at $p$ with the corresponding $\limsup\limits_{x\rightarrow\infty}\frac{\log|f_{i}(x)|}{\log d(x,p)}$ arranged in a non-decreasing order for $1\leq i\leq n$, we denote this set by $\mathcal{O}_{P}(M,p)$.
+In other words, for any $1\leq k\leq n$ we have $d_{\operatorname{min}}^{(k)}=\inf\limits_{f_{k}}\limsup\limits_{x\rightarrow\infty}\frac{\log|f_{k}(x)|}{\log d(x,p)}$, where the infimum is taken among all possible $f_{k}$ that appears in the $k$-th component of some sequence in $\mathcal{O}_{P}(M,p)$.
+Note that apriori it’s unclear if $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$ can be obtained by an n-tuple holomorphic functions in $\mathcal{O}_{P}(M,p)$.
 But obviously $d_{\operatorname{min}}=d_{\operatorname{min}}^{(1)}$.
-Then Yang proposed the following conjecture in [27] on the relation between the above quantities which can be understood as the quantitative version of Theorem 1.2.
-Conjecture 1.5 (Conjecture 2.5.6, Conjecture 2.5.8 in [27]).
+Then Yang proposed the following conjecture on the relation between the above quantities which can be understood as the quantitative version of Theorem 1.
+Conjecture 2 (Conjecture 2.5.6, Conjecture 2.5.8 in [Yang13]).
 Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature. Assume the universal cover of $M$ does not split. Then
-(1) $\overrightarrow{d_{\operatorname{min}}}(p)$ can be realized by a n-tuple holomorphic functions in $\mathcal{O}_{P,p}(M)$ and is independent of the choice of $p$.
+(1) If $M$ is of maximal volume growth, then $\overrightarrow{d_{\operatorname{min}}}(p)$ can be realized by a n-tuple holomorphic functions in $\mathcal{O}_{P}(M,p)$ and is independent of the choice of $p$.
 (2) $\operatorname{AVR}(M,g)=\prod\limits_{i=1}^{n}\frac{1}{d_{\operatorname{min}}^{(i)}}$.
@@
 (4) $\operatorname{ASCD}(M,g)=4nD_{\operatorname{min}}$.
-In the case of nonmaximal volume growth, (1),(2) are true because $d_{\min}=+\infty$ from Theorem 1.2, and $\operatorname{ASCD}=+\infty$.
-Assume $M$ is of maximal volume growth, in [27], Yang verified the conjecture is true for $U(n)$-invariant Kähler metrics on $\mathbb{C}^{n}$.
-Liu proved (1) implicitly in [16].
+In the case of nonmaximal volume growth, (2) is true because $d_{\min}=+\infty$ from Theorem 1. Also we can show that $\operatorname{ASCD}=+\infty$.
+Assume $M$ is of maximal volume growth, in [Yang13], Yang verified the conjecture is true for $U(n)$-invariant Kähler metrics on $\mathbb{C}^{n}$.
+Liu proved (1) implicitly in [Liu19].
 In fact, he showed that any n-tuple of polynomial growth functions which are algebraically independent and of minimal degrees can form a global coordinate.
-Then recently in [18], Liu proved (2) and derived the explicit formula $\operatorname{ASCD}(M,g)=4n(\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n)$ by passing the geometric quantities to the tangent cones and exploiting the metric Kähler cone structure of the limit space.
-In [28], Yang proved that (3) holds for gradient expanding Kähler-Ricci solitons with nonnegative Ricci curvature by using the Poincaré coordinate introduced in [1], see Theorem 1.8.
-Therefore, the general proof of (3) remained the final step toward resolving Conjecture 1.5 in full generality.
-In this paper, we prove that (3) of Conjecture 1.5 is true, thereby completely resolving the conjecture. Precisely, we proved
+Then recently in [Liu24], Liu proved (2) and derived the explicit formula $\operatorname{ASCD}(M,g)=4n(\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n)$ by passing the geometric quantities to the tangent cones and exploiting the metric Kähler cone structure of the limit space.
+In [Yang22], Yang proved that (3) holds for gradient expanding Kähler-Ricci solitons with nonnegative Ricci curvature by analysing the Poincaré coordinate introduced in [Bry08], see Theorem 3 for more details.
+Therefore, the general proof of (3) remained the final step toward resolving Conjecture 2 in full generality.
+In this paper, we prove that (3) of Conjecture 2 is true, thereby completely resolving the conjecture. Precisely,
 Theorem A.
@@
 Now we move to the behavior of the Kähler-Ricci flow on complete noncompact Kähler manifolds where $(M,g)$ has nonnegative bisectional curvature.
-$\frac{\partial g(t)}{\partial t}=-\operatorname{Ric}(g(t)),\quad g(0)=g.$ (1.1)
-When the initial metric $(M,g)$ has bounded curvature and Euclidean volume growth, the long-time behavior of (1.1) has been described very clearly. Let us summarize these results in the following:
-Theorem 1.6 (Theorem 1.2, Proposition 3.2 in [3], Theorem 6.1 in [4]).
-Suppose that $(M^{n},J,g)$ is a complete noncompact Kähler manifold, where $g$ has bounded nonnegative bisectional curvature and Euclidean volume growth with $\operatorname{AVR}(M,g)=\nu>0$.
+(1.1) $\frac{\partial g(t)}{\partial t}=-\operatorname{Ric}(g(t)),\quad g(0)=g.$
+When the initial metric $(M,g)$ has maximal volume growth, the long-time behavior of (1.1) has been described very clearly. Let us summarize these results in the following:
+Theorem 2 (Corollary 1 in [Ni05], Theorem 1.2, Proposition 3.2 in [CT06], Theorem 6.1 in [CT11], Theorem 1.5 in [LT20], Theorem 1.1 in [Lee25]).
+Suppose that $(M^{n},J,g)$ is a complete noncompact Kähler manifold, where $g$ has nonnegative bisectional curvature and maximal volume growth with $\operatorname{AVR}(M,g)=\nu>0$.
 Then the following conclusions hold:
-(1) The solution to the Kähler-Ricci flow with the initial metric $g$ exists for all $t\in[0,+\infty)$ and it has nonnegative bisectional curvature for any $t\geq 0$.
-Moreover, there exists a constant $C(n,\nu)$ such that $\|\operatorname{Rm}(x,t)\|\leq\frac{C(n,\nu)}{t}$ for any $x\in M$ and $t\in(0,+\infty)$.
+(1) There exists a unique long-time solution $g(t)$ to the Kähler-Ricci flow on $M\times[0,+\infty)$ with the initial metric $g$ such that,
+  1. (i)
+    It has nonnegative bisectional curvature for any $t\geq 0$.
+  2. (ii)
+    There exists a constant $C(n,\nu)$ such that $\|\operatorname{Rm}(x,t)\|\leq\frac{C(n,\nu)}{t}$ for any $x\in M$ and $t\in(0,+\infty)$.
+  3. (iii)
+    $\operatorname{AVR}(M,g(t))=\operatorname{AVR}(M,g)=\nu$.
 (2) For any point $x\in M$, let $\{\lambda_{1}(x,t),\cdots,\lambda_{n}(x,t)\}$ denote the eigenvalues of Ricci curvature $\operatorname{Ric}(x,t)$ with respect to $g(t)$ in the nondecreasing order.
-Then $t\lambda_{i}(x,t)$ is nondecreasing on $t>0$, hence $\mu_{i}(x):=\lim_{t\rightarrow+\infty}t\lambda_{i}(x,t)$ exists.
+Then $t\lambda_{i}(x,t)$ is nondecreasing on $t>0$, hence $\mu_{i}(x):=\lim\limits_{t\rightarrow+\infty}t\lambda_{i}(x,t)$ exists.
 (3) If $\mu_{1}(x)<\mu_{2}(x)<\cdots<\mu_{l}(x)$ are the distinct limits in (2), where $l\leq n$.
@@
 Moreover, both convergences are uniform over all $v\in V_{i}\backslash\{0\}$.
-$\sum\limits_{i=1}^{l}-\mu_{i}(x)\operatorname{dim}_{\mathbb{C}}V_{i}=\lim\limits_{t\to\infty}\log\frac{\det(g_{i\bar{j}}(x,t))}{\det(g_{i\bar{j}}(x,0))}.$ (1.2)
+(1.2) $\sum\limits_{i=1}^{l}-\mu_{i}(x)\operatorname{dim}_{\mathbb{C}}V_{i}=\lim\limits_{t\to\infty}\log\frac{\det(g_{i\bar{j}}(x,t))}{\det(g_{i\bar{j}}(x,0))}.$
 (5) Fix any point $p\in M$.
@@
     $(N,J_{\infty},h(t))$ has nonnegative bisectional curvature for any $t>0$, and there exists a real-valued function $f\in C^{\infty}(N)$ such that $(N,J_{\infty},h(1))$ satisfies the expanding soliton equation
-    $\operatorname{Ric}_{i\bar{j}}(h(1))+h_{i\bar{j}}(1)-f_{i\bar{j}}=0,\quad f_{ij}=f_{\bar{i}\bar{j}}=0.$ (1.3)
+    (1.3) $\operatorname{Ric}_{i\bar{j}}(h(1))+h_{i\bar{j}}(1)-f_{i\bar{j}}=0,\quad f_{ij}=f_{\bar{i}\bar{j}}=0.$
     Moreover, $\nabla_{h(1)}f(O)=0$ and the eigenvalues of Ricci curvature of $h(1)$ at $O$, arranged in the non-decreasing order, equal $\mu_{i}(p)$ for $1\leq i\leq n$.
-Indeed, in (3) in Theorem 1.6, $\mu_{i}$ play the role of Lyapunov exponents in a dynamical-system interpretation on asymptotics of (1.1), it basically says that the Ricci curvature can be “simultaneously diagonalized” near $t=+\infty$ in some sense.
-See Theorem 4.1 in [3] for a precise statement.
-Then Yang proposed the following conjecture in [27]:
-Conjecture 1.7 (Conjecture 2.5.16 in [27]).
-Let $(M^{n},g)$ be a complete noncompact Kähler manifold with bounded nonnegative curvature and Euclidean volume growth.
-Let $g(t)$, $t\in[0,+\infty)$, be the complete solution Kähler-Ricci flow with initial metric $g$, then
+Remark.
+(1) If $(M,g)$ has bounded curvature, the above theorem was proved by Chau–Tam in [CT06, CT11] through Li-Yau-Hamilton estimate.
+In the case of unbounded curvature, the existence of such solution was established by Lee–Tam in [LT20, LT21], in which they constructed a pyramid solution to the Kähler–Ricci flow.
+Moreover, the uniqueness of complete non-compact Ricci flows under a scaling-invariant curvature bound was proved very recently by Lee [Lee25].
+(2) In the context of the Kähler–Ricci flow, the theorem asserts that $(M,J,g(t))$ asymptotically approaches an expanding gradient Kähler–Ricci soliton at the base point $p$.
+Moreover, conclusions (2) and (3) essentially indicate that the Ricci curvature $\operatorname{Ric}(x,t)$ becomes “simultaneously diagonalizable” as $t\rightarrow+\infty$, in an appropriate asymptotic sense.
+In fact, $\mu_{i}$’s play the role of Lyapunov exponents in a dynamical-system interpretation on asymptotics of (1.1).
+Then Yang proposed the following conjecture in [Yang13]:
+Conjecture 3 (Conjecture 2.5.16 in [Yang13]).
+Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative curvature and maximal volume growth.
+Let $g(t)$, $t\in[0,+\infty)$, be the unique complete solution Kähler-Ricci flow with initial metric $g$ in Theorem 2, then
 (1) $\mu_{i}(p)$ is independent of the choice of $p$ for any $1\leq i\leq n$.
@@
 (3) $\operatorname{ASCD}(M,g(t))$ is invariant along $g(t)$.
-In [28], Yang proved (1) using Shi’s curvature estimates. He further proved (2) for expanding Kähler-Ricci solitons with nonnegative Ricci curvature. Precisely,
-Theorem 1.8 (Theorem 1.4 in [28]).
+In [Yang22], Yang proved (1) using Shi’s curvature estimates on Ricci flow solutions under a scaling-invariant curvature bound.
+He further proved (2) for expanding Kähler-Ricci solitons with nonnegative Ricci curvature.
+Precisely,
+Theorem 3 (Theorem 1.4 in [Yang22]).
 Let $(N^{n},J,O,g,f)$ be a complete noncompact gradient Kähler-Ricci soliton with nonnegative Ricci curvature, normalized so that $f=R+|\nabla^{g}f|^{2}$.
-Let $\mu_{1}\leq\cdots\leq\mu_{n}$ be the eigenvalues of Ricci curvature at $O$.
+Let $0\leq\mu_{1}\leq\cdots\leq\mu_{n}$ be the eigenvalues of Ricci curvature at $O$.
 Then
-$\overrightarrow{d_{\operatorname{min}}}(q)=\{\mu_{1}+1,\cdots,\mu_{n}+1\}\text{ for all }q\in N;$ (1.4)
+(1.4) $\overrightarrow{d_{\operatorname{min}}}(q)=\{\mu_{1}+1,\cdots,\mu_{n}+1\}\text{ for all }q\in N;$
 and
-$D_{\operatorname{min}}=\sum\limits_{i=1}^{n}\mu_{i}.$ (1.5)
-If $(M,g)$ has unbounded curvature, in [8], Lee-Tam established the following long-time existence result for the Kähler-Ricci flow:
-Theorem 1.9 (Corollary 1 in [21], Theorem 1.5 in [8]).
-Suppose $(M^{n},g)$ is a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth, $\operatorname{AVR}(M,g)=\nu>0$.
-Then there exists $C=C(n,\nu)>0$ depending only on $n,\nu$ such that there exists a smooth Kähler-Ricci flow solution $g(t)$ on $[0,+\infty)$ such that
-$\|\operatorname{Rm}\|(g(t))\leq\frac{C(n,\nu)}{t}$
-on $M\times(0,+\infty)$.
-(2) $g(t)$ has nonnegative bisectional curvature.
-(3) $\operatorname{AVR}(g(t))=\operatorname{AVR}(g(0))=\nu$ on $M\times(0,+\infty)$.
-We emphasize that the uniqueness of such Kähler-Ricci flow solution remains unknown in this setting.
-Consequently, though these $\mu_{i}$’s can be defined for $M$ with unbounded curvature, a priori, we cannot guarantee that they are independent of the particular flow solution g(t).
-In this paper, we prove the Conjecture 1.7 holds in general.
+(1.5) $D_{\operatorname{min}}=\sum\limits_{i=1}^{n}\mu_{i}.$
+In this paper, we show that the Conjecture 3 holds in general.
 Theorem B.
-Conjecture 1.7 holds.
-Since Conjecture 1.7 is true, then $\mu_{i}$’s are indeed independent of $g(t)$ and thus well-defined for manifolds with unbounded curvature.
-Therefore we can remove the boundedness condition in Conjecture 1.7.
-Corollary 1.10.
-Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative curvature and Euclidean volume growth.
-Let $g(t)$ $(t\in[0,+\infty)$ be as complete solution Kähler-Ricci flow with initial metric $g$ as in Theorem 1.9, then
-(1) $\mu_{i}(p)$ is independent of the choice of $p$ for any $1\leq i\leq n$.
-(2) $\mu_{i}=d_{\operatorname{min}}^{(i)}-1$ for any $1\leq i\leq n$.
-(3) $\operatorname{ACSD}(M,g(t))$ is invariant along $g(t)$.
-Note that by Claim 4.2 in [16], the degrees of functions in $\mathcal{O}_{P}(M)$ also can be “simultaneously diagonalized” along the blowdown sequence, for details see Theorem 2.4 (5) in Section 2.
-As a by-product of Corollary 1.10, we can describe the Lyapunov regularity in (3) of Theorem 1.6 using the polynomial ring $\mathcal{O}_{P}(M)$.
+Conjecture 3 holds.
+Note that by Claim 4.2 in [Liu19], the degrees of polynomial growth functions also can be “simultaneously diagonalized” along the metric blow-down sequence, we will discuss the phenomenon in Theorem 5 (5) in Section 2.
+As a by-product, we can describe the Lyapunov regularity in (3) of Theorem 2 using the polynomial ring $\mathcal{O}_{P}(M)$.
 In some sense, it bridges the two distinct proofs of Yau’s uniformization conjecture in the case of maximal volume growth.
-Corollary 1.11.
-In (3) of Theorem 1.6, the basis of $V_{i}$ can be taken as $\left\{\frac{\partial}{\partial f_{i1}}(x),\cdots,\frac{\partial}{\partial f_{im_{i}}}(x)\right\}$.
+Corollary 1.
+In (3) of Theorem 2, the basis of $V_{i}$ can be taken as $\left\{\frac{\partial}{\partial f_{i1}}(x),\cdots,\frac{\partial}{\partial f_{im_{i}}}(x)\right\}$.
 Here $\operatorname{dim}V_{i}=m_{i}$ is the multiplicity of $\mu_{i}$, $\sum\limits_{i=1}^{l}m_{i}=n$, and
@@
 for any $1\leq s\leq m_{i}$.
-Moreover, $\{f_{11},\cdots,f_{1m_{1}},\cdots,f_{l1},\cdots,f_{lm_{l}}\}$ can serve as a biholomorphism from $M$ onto $\mathbb{C}^{n}$.
+Moreover, $\{f_{11},\cdots,f_{1m_{1}};\cdots;f_{l1},\cdots,f_{lm_{l}}\}$ can serve as a biholomorphism from $M$ onto $\mathbb{C}^{n}$.
+Now let’s turn our attention to the tangent cone of $(M,g)$ with nonnegative bisectional curvature and maximal volume growth.
+By Lott [Lott21], any tangent cone of $M$ is a smooth complex manifold which is biholomorphic to $\mathbb{C}^{n}$ with Kähler current $\omega=\frac{\sqrt{-1}}{2}\partial\bar{\partial}r^{2}$, where $r$ is the radial function on the tangent cone.
+Moreover, it satisfies the condition $\operatorname{BK}\geq 0$”, which is a new pluripotential-theoretic characterization which is equivalent to nonnegative bisectional curvature on Kähler manifolds, and can pass to Gromov-Hausdorff limits.
+In this paper, we shall prove the following structure theorem about the tangent cones:
+Theorem C.
+Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth.
+Let $(M_{\infty}=C(X),d)$ be a tangent cone at infinity of $M$, which is biholomorphic to $\mathbb{C}^{n}$, with link $X$.
+(1) $X$ is homeomorphic to the sphere $\mathbb{S}^{2n-1}$.
+(2) If we assume $(M,g)$ has quadratic curvature decay, then there exists a sequence of smooth Kähler cone metric $\omega_{i}=\frac{\sqrt{-1}}{2}\partial\bar{\partial}r_{i}^{2}$, which have the same Reeb field with $M_{\infty}$, such that $(\mathbb{C}^{n},\omega_{i})$ converges to $(M_{\infty},d)$ in the Gromov-Hausdorff sense.
+In particular, for any fixed point $p\in\mathbb{C}^{n}$, $\operatorname{BK}(\omega_{i},p)\geq-\Psi(\frac{1}{i})$.
+Here, and below, $\Psi(\frac{1}{i})$ denotes a positive function converging to zero as $i\rightarrow+\infty$.
+Remark.
+(1) Theorem C (1) can be seen as a Kähler analogue of Theorem 1.3 in Kapovitch [Kap02], which states that the link of any tangent cone of a complete noncompact Riemannian manifold with nonnegative sectional curvature and maximal volume growth has Alexandrov curvature bounded below by $1$ and is homeomorphic to a sphere.
+(2) From Theorem 1 we know that if $M$ is of maximal volume growth, then the curvature is of average quadratic decay.
+So the assumption of (2) doesn’t seem to be that strict.
+Nevertheless, we still wonder if this condition can be removed.
+Also, we wonder if we can get a uniform bisectional curvature lower bound, i.e.
+$\operatorname{BK}(\omega_{i})\geq-\frac{\Psi(\frac{1}{i})}{r_{i}^{2}}$ uniformly.
 This paper is organized as follows.
 Section 2 is some basic preliminary results and some simple conclusions that will be used later.
-In section 3, we will prove Theorem B by noting that any tangent cone of $M$ also arises as a tangent cone of corresponding expanding Kähler-Ricci solitons in Theorem 1.6.
-Also we prove Corollary 1.11.
+In Section 3, we will prove Theorem B by noting that any tangent cone of $M$ also arises as a tangent cone of corresponding expanding Kähler-Ricci solitons described in Theorem 2.
+After finishing the proof, we will prove Corollary 1.
 Section 4 contains the proof of Theorem A.
-Acknowledgements
-I would like to express my sincere gratitude to my advisor, Professor Yihu Yang, for providing constant support. I would like to extend my thanks to Professor Gang Liu for many valuable discussions.
+In Section 5 we will prove Theorem C using some knowledge of Sasaki geometry and discuss several further questions.
+Acknowledgement: I would like to express my sincere gratitude to my advisor, Yihu Yang, for his constant support and lots of inspiring discussions.
+I would like to extend my thanks to Gang Liu and Zhiqin Lu for many valuable discussions.
+I am thankful to Man-Chun Lee for his interest in this work and valuable comments.
+I am grateful to University of California, Irvine for hospitality during my visit in fall 2025.
 2. Preliminary results
-2.1. Structure of Kähler manifolds with maximal volume growth and the tangent cones
-We first recall the main result in [16], and prove some relevant facts that will be used later.
-Theorem 2.1.
-Let $(M^{n},g,p)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth. Let $\operatorname{AVR}(M,g)=\nu>0$. Then
+2.1. Structure of Kähler manifolds with maximal volume growth and tangent cones
+We first collect several results about the strucutures of complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth and their tangent cones, and prove some relevant facts that will be used later.
+The following is the main theorem in [Liu19].
+Theorem 4.
+Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth, $p\in M$. Let $\operatorname{AVR}(M,g)=\nu>0$. Then
 (1) $M$ is biholomorphic to $\mathbb{C}^{n}$.
 (2) There exists an n-tuple polynomial growth holomorphic functions $(f_{1},\cdots,f_{n})$ serving as a proper biholomorphism onto $\mathbb{C}^{n}$. These functions satisfy
-$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p,1)}f_{i}\overline{f_{j}}=\delta_{ij},\quad f_{i}(p)=0\quad\text{ for any }i,j\in\{1,\cdots,n\}.$ (2.1)
+(2.6) $\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p,1)}f_{i}\overline{f_{j}}=\delta_{ij},\quad f_{i}(p)=0\quad\text{ for any }i,j\in\{1,\cdots,n\}.$
 Moreover, any element in $\mathcal{O}_{P}(M)$ is a polynomial of these functions, meaning that $\mathcal{O}_{P}(M)\cong\mathbb{C}[f_{1},\cdots,f_{n}]$.
@@
 Proof.
-The proof of these conclusions is implicitly contained in [11, 16]. For reader’s convenience, we give a thorough explanation. (1) and (2) are contained in Section 4 in [16].
-The properness is also contained in [11].
-It was shown that there exists a constant $D=D(n,\nu)$ such that let $(g_{1},\cdots,g_{k})$ be a linearly independent basis of $\mathcal{O}_{P}(M)$, where $k=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{P}(M)-1$, such that
-$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p,1)}g_{i}\overline{g_{j}}=\delta_{ij},\quad g_{i}(p)=0\quad\text{ for any }i,j\in\{1,\cdots,k\}.$
+The proof of these conclusions is implicitly contained in [Liu16a, Liu19]. For reader’s convenience, we give a thorough explanation. The proof of (1) and (2) is contained in Section 4 in [Liu19].
+The properness is also contained in [Liu16a].
+It’s shown that there exists a constant $D=D(n,\nu)$ such that let $(g_{1},\cdots,g_{k})$ be a linearly independent basis of $\mathcal{O}_{D}(M)$, where $k=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{D}(M)-1$, such that
+$\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p,1)}g_{i}\overline{g_{j}}=\delta_{ij},\quad g_{i}(p)=0\quad\text{ for any }i,j\in\{1,\cdots,k\}.$
 and
@@
 Take an $n$-tuple algebraically independent holomorphic functions with minimal degrees in $(g_{1},\cdots,g_{k})$, denoted by $(f_{1},\cdots,f_{n})$.
-As in section 4 in [16], $(f_{1},\cdots,f_{n})$ serves as a biholomorphism of $M$ onto $\mathbb{C}^{n}$, and any polynomial growth function is a polynomial of them.
-Since any $g_{i}$ is a polynomial of these $(f_{1},\cdots,f_{n})$ with degree no larger than $D$ by the three-circle theorem, by a contradiction argument we get
-$\min\limits_{\partial B(p,r)}\sum\limits_{i=1}^{n}|f_{i}|^{2}\geq c(n,\nu)r^{\frac{2}{D}}$ (2.2)
+Following the proof in section 4 in [Liu19], $(f_{1},\cdots,f_{n})$ serves as a biholomorphism of $M$ onto $\mathbb{C}^{n}$, and any polynomial growth function is a polynomial of them.
+Since any $g_{i}$ is a polynomial of these $(f_{1},\cdots,f_{n})$ with degree at most $D$ by the three-circle theorem, by a contradiction argument we get
+(2.7) $\min\limits_{\partial B(p,r)}\sum\limits_{i=1}^{n}|f_{i}|^{2}\geq c(n,\nu)r^{\frac{2}{D}}$
 holds for any $r>0$ and some constant $c=c(n,\nu)$. In other words, $(f_{1},\cdots,f_{n})$ is proper.
 For (3), given any two such global coordinates $(f_{1},\cdots,f_{n})$ and $(h_{1},\cdots,h_{n})$.
-Assume their degrees are arranged in the nondecreasing order, then the matrix $\left(\frac{\partial h_{i}}{\partial f_{j}}\right)$ and $\left(\frac{\partial f_{i}}{\partial h_{j}}\right)$ are both polynomial matrices of $(f_{1},\cdots,f_{n})$ and are inverse of each other.
+Assume their degrees are both arranged in the nondecreasing order, then the matrix $\left(\frac{\partial h_{i}}{\partial f_{j}}\right)$ and $\left(\frac{\partial f_{i}}{\partial h_{j}}\right)$ are both polynomial matrices of $(f_{1},\cdots,f_{n})$ and are inverse of each other.
 Therefore they must both be constant matrices.
@@
 ∎
-In fact, the minimal degrees of all coordinates in Theorem 2.1 (3) described above are just the refined minimal order $\overrightarrow{d_{\operatorname{min}}}$ defined in the introduction, and it has been shown in [16].
-Corollary 2.2.
-Conjecture 1.5 (1) holds.
+In fact, the minimal degrees of the coordinate in Theorem 4 (3) described above are just the refined minimal order $\overrightarrow{d_{\operatorname{min}}}$ defined in the introduction, and it has been shown in [Liu19].
+Corollary 2.
+Conjecture 2 (1) holds.
 Proof.
 If $M$ is of nonmaximal volume growth.
-By Theorem 2 in [12], $\mathcal{O}_{P}(M)=\mathbb{C}$, i.e.
+By Theorem 2 in [Liu16b], $\mathcal{O}_{P}(M)=\mathbb{C}$, i.e.
 $\overrightarrow{d_{\operatorname{min}}}=\overrightarrow{+\infty}$, which is independent of the choice of $p$.
-If $M$ is of maximal volume growth, due to [16], there exists a strictly increasing sequence $1\leq d_{1}<d_{2}<\cdots$, so that $\mathcal{O}_{P}(M)=\mathbb{C}\oplus\left(\mathcal{O}_{d_{1}}(M)/\mathbb{C}\right)\oplus\left(\mathcal{O}_{d_{2}}(M)/\mathcal{O}_{d_{1}}(M)\right)\oplus\cdots$ as a complex vector space.
+If $M$ is of maximal volume growth, due to [Liu19], there exists a strictly increasing sequence $1\leq d_{1}<d_{2}<\cdots$, so that $\mathcal{O}_{P}(M)=\mathbb{C}\oplus\left(\mathcal{O}_{d_{1}}(M)/\mathbb{C}\right)\oplus\left(\mathcal{O}_{d_{2}}(M)/\mathcal{O}_{d_{1}}(M)\right)\oplus\cdots$ as a complex vector space.
 For any $k\in\mathbb{N}$, pick a maximal linearly independent vectors $f_{k1},\cdots,f_{km_{k}}$ of $\mathcal{O}_{d_{k}}(M)$ so that they form a basis of $\mathcal{O}_{d_{k}}(M)/\mathcal{O}_{d_{k-1}}(M)$ as quotient of vector spaces and no element in the span is given by polynomials of $\mathcal{O}_{d_{k-1}}(M)$.
-Then the first $n$ functions $\{f_{11},\cdots,f_{1m_{1}},f_{21},\cdots\}$ form a global coordinate in Theorem 2.1 after normalization as (2.1).
+Then the first $n$ functions $\{f_{11},\cdots,f_{1m_{1}},f_{21},\cdots\}$ form a global coordinate in Theorem 4 after normalization as (2.6).
 […]
-Remark 2.3.
-(1) As can be seen in the proof, if $M$ is of maximal volume growth, Corollary 2.2 still holds without the nonsplitting condition on the universal cover.
-(2) For simplicity, if $M$ is of maximal volume growth we just call the coordinate as in (2) in Theorem 2.1 a canonical coordinate on $M$, with degrees $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$.
+Remark.
+(1) As can be seen in the proof, if $M$ is of maximal volume growth, Corollary 2 still holds without the nonsplitting condition on the universal cover.
+(2) For simplicity, if $M$ is of maximal volume growth we just call the coordinate as in (2) in Theorem 4 a canonical coordinate on $M$, with degrees $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$.
 And we arrange them in non-decreasing order of their degrees.
-Let $M_{\infty}$ be a tangent cone of infinity of $M$, whose structure has been studied in [14, 15, 9, 19, 18]. Let’s summarize these results in the following:
-Theorem 2.4.
+Now we recall the definition of $\operatorname{BK}\geq K$” in Lott [Lott21].
+Definition 2.
+Let $X$ be a reduced complex space, then for each $x\in X$, there is a neighborhood $U_{x}$ of $x$ and an embedding $e_{x}:U_{x}\rightarrow\mathbb{C}^{N_{x}}$ so that $e\left(U_{x}\right)$ is the zero set of a finite number of analytic functions defined on an open set $V_{x}\subset\mathbb{C}^{N_{x}}$.
+(1) We call $X$ a Kähler space if on each $U_{x}$, there exists a continuous plurisubharmonic function $\phi_{x}$, so that $\phi_{x}-\phi_{y}$ is pluriharmonic on each $U_{x}\cap U_{y}\neq\varnothing$.
+Thus $\omega=\sqrt{-1}\partial\bar{\partial}\phi_{x}$ on $U_{x}$ is a well-defined positive (1,1)-current, and we call it the Kähler current on $X$.
+(2) Let $d$ be a metric on a Kähler space $(X,\omega)$, we call $(X,\omega,d)$ a metric Kähler space, if the metric $d$ induces the topology of $X$, so that if $\Sigma$ is an embedded holomorphic disk, then for all $x\in X$, $\omega|_{\Sigma}$ equals the 2-dimensional Hausdorff measure on each $\Sigma\cap U_{x}\neq\varnothing$.
+(3) Given $(X,\omega,d)$ a metric Kähler space and $K\in\mathbb{R}$, let $d_{p}\in C(X)$ be the distance function from $p\in X$. Define $d_{p,K}\geq 0$ by
+$d_{p,K}^{2}=\begin{cases}-\frac{4}{K}\log\cos\left(d_{p}\sqrt{\frac{K}{2}}\right)&\text{ if }K>0,\\ d_{p}^{2}&\text{ if }K=0,\\ \frac{4}{-K}\log\cosh\left(d_{p}\sqrt{\frac{-K}{2}}\right)&\text{ if }K<0.\end{cases}$
+(If $K>0$, then we restrict to $d_{p}\leq\frac{\pi}{\sqrt{2K}}$.)
+Then we call $X$ has $\operatorname{BK}\geq K$” if for any $x\in X$, $\phi_{x}-d_{p,K}^{2}/2$ is plurisubharmonic on $U_{x}$ for any $p\in X$.
+Moreover, if $X$ is a smooth Kähler manifold, the condition $\operatorname{BK}\geq K$” is equivalent to that bisectional curvature has lower bound $K$.
+The structure of tangent cones of infinity of $M$ has been studied in [Liu18a, Liu18b, LT21, Lott21, Liu24]. Let’s summarize these results in the following:
+Theorem 5.
 Let $(M,g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth.
@@
 (1) $(M_{\infty},p_{\infty},d_{\infty})$ is a complex manifold which is biholomorphic to $\mathbb{C}^{n}$.
-Moreover, it’s a metric Kähler cone with $BK\geq 0$” (see Section 5 in [19] for relevant definitions).
-(2) There exists a complete Kähler-Ricci flow solution $h(t)$ on $M_{\infty}$ for $t\in[0,+\infty)$ with nonnegative bisectional curvature such that $(M_{\infty},p_{\infty},d_{\infty})$ is also the Gromov-Hausdorff limit of $h(t)$ as $t\rightarrow 0$.
-Moreover, $(M_{\infty},h(t))$ is a gradient Kähler-Ricci soliton with nonnegative bisectional curvature.
+Moreover, it has $BK\geq 0$” with Kähler current $\omega=\frac{\sqrt{-1}}{2}\partial\bar{\partial}r^{2}$, where $r$ is the radial function on $M_{\infty}$.
+(2) There exists a complete family of gradient expanding Kähler-Ricci soliton coming out of $M_{\infty}$, i.e.
+there exists $(M_{\infty},h(t))$ a family of gradient expanding Kähler-Ricci soliton with nonnegative bisectional curvature such that $h(t)\rightarrow d_{\infty}$ as $t\rightarrow 0$ in the Gromov-Hausdorff sense.
 (3) $\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M_{\infty})=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)$ for any $d\geq 0$.
 And there exists a n-tuple of polynomial growth functions $(z_{1\infty},\cdots,z_{n\infty})$ serving as a proper biholomorphism onto $\mathbb{C}^{n}$.
-Each $z_{i\infty}$ is of homogeneous degree $d_{\operatorname{min}}^{(i)}$, i.e.
-there exists a eigenfunction $\phi_{i}$ of the Laplacian operator on the link such that $z_{i\infty}=r^{d_{\operatorname{min}}^{(i)}}\phi_{i}$ on $M_{\infty}$.
+Each $z_{i\infty}$ is of homogeneous degree $d_{\operatorname{min}}^{(i)}$.
 They satisfy
-$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p_{\infty},1)}z_{i\infty}\overline{z_{j\infty}}=\delta_{ij},\quad z_{i\infty}(p_{\infty})=0\quad\text{ for any }i,j\in\{1,\cdots,n\}.$ (2.3)
+(2.8) $\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p_{\infty},1)}z_{i\infty}\overline{z_{j\infty}}=\delta_{ij},\quad z_{i\infty}(p_{\infty})=0\quad\text{ for any }i,j\in\{1,\cdots,n\}.$
 Moreover, any element in $\mathcal{O}_{P}(M_{\infty})$ is a polynomial of these functions. In particular, $\mathcal{O}_{P}(M_{\infty})\cong\mathbb{C}[z_{1\infty},\cdots,z_{n\infty}]$.
@@
 Normalize them to $(z_{1k},\cdots,z_{nk})$ on $M_{k}$ such that
-$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p_{k},1)}z_{ik}\overline{z_{jk}}=\delta_{ij},\quad z_{ik}(p_{k})=0\quad\text{ for any }i,j\in\{1,\cdots,n\},k\in\mathbb{N}.$ (2.4)
+(2.9) $\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p_{k},1)}z_{ik}\overline{z_{jk}}=\delta_{ij},\quad z_{ik}(p_{k})=0\quad\text{ for any }i,j\in\{1,\cdots,n\},k\in\mathbb{N}.$
 Then $(z_{1k},\cdots,z_{1n})$ sub-sequentially converges to a global coordinate $(z_{1\infty},\cdots,z_{n\infty})$ as in (3) on $M_{\infty}$ uniformly on compact subsets as $k\rightarrow+\infty$.
-(6) The volume of the unit ball in $M_{\infty}$ is $\frac{\omega_{2n}}{d_{\operatorname{min}}^{(1)}\cdots d_{\operatorname{min}}^{(n)}}$. In particular, Conjecture 1.5 (2) holds.
+(6) The volume of the unit ball in $M_{\infty}$ is $\frac{\omega_{2n}}{d_{\operatorname{min}}^{(1)}\cdots d_{\operatorname{min}}^{(n)}}$. In particular, Conjecture 2 (2) holds.
 Proof.
-(1) and (2) were proved in Proposition 6.1 in [19].
-For (3), for any $d\geq 0$, $\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M_{\infty})=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)$ by Proposition 4.1 in [16].
-Now we show the existence of such coordinate.
+(1) and (2) are proved in Proposition 6.1 in [Lott21].
+For (3), for any $d\geq 0$, $\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M_{\infty})=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)$ by Proposition 4.1 in [Liu19].
+Next we show the existence of such coordinate.
 Firstly we claim that the homothetic vector field $r\frac{\partial}{\partial r}$ is holomorphic on $M_{\infty}$.
-By Theorem 1.9 there exists a long-time Kähler-Ricci flow solution $g(t)$ on $M$ such that $(M_{\infty},J_{\infty},p_{\infty},h(t))$ is the pointed Cheeger-Hamilton limit of a blow-down sequence of $g(t)$, i.e.
-$(M,J,p,g_{i}(t)=\frac{1}{t_{i}}g(t_{i}t)$ for some sequence $t_{i}\rightarrow+\infty$.
+By Theorem 2 (1), (5) there exists a long-time Kähler-Ricci flow solution $g(t)$ on $M$ such that $(M_{\infty},J_{\infty},p_{\infty},h(t))$ is the pointed Cheeger-Hamilton limit of a blow-down sequence of $g(t)$, i.e.
+$(M,J,p,g_{i}(t)=\frac{1}{t_{i}}g(t_{i}t))$ for some sequence $t_{i}\rightarrow+\infty$.
 In other words, $J$ converges to $J_{\infty}$ through a sequence of diffeomorphisms exhausted on $M_{\infty}$.
-By Claim 4.3 in [16], the complexification of $r\frac{\partial}{\partial r}$ is in the span of several holomorphic vector fields on $M_{\infty}$.
+By Claim 4.3 in [Liu19], the complexification of $r\frac{\partial}{\partial r}$ is in the span of several holomorphic vector fields on $M_{\infty}$.
 Moreover, $(r\frac{\partial}{\partial r})\mathcal{O}_{d}(M_{\infty})\subset\mathcal{O}_{d}(M_{\infty})$ holds for any $d\geq 0$.
 The claim is confirmed.
-Therefore, $r\frac{\partial}{\partial r}$ is a contracted holomorphic field on $M_{\infty}$, from a result in [25], $M_{\infty}$ is biholomorphic to $\mathbb{C}^{n}$.
-Now follow Section 4 in [16], replace the holomorphic vector field $X$ in the proof by $r\frac{\partial}{\partial r}$, we conclude the proof of (3) except the homogenity of these coordinates.
+Therefore, $r\frac{\partial}{\partial r}$ is a contracted holomorphic field on $M_{\infty}$, from a result in [RR88], $M_{\infty}$ is biholomorphic to $\mathbb{C}^{n}$.
+Now follow Section 4 in [Liu19], replace the holomorphic vector field $X$ in the proof by $r\frac{\partial}{\partial r}$, we conclude the proof of (3) except the homogenity of these coordinate functions.
 This is also obvious since we can just choose the highest order terms of each coordinate functions and they certainly form a new biholomorphism onto $\mathbb{C}^{n}$.
-The proof of (4) is the same as the proof of (3) in Theorem 2.1.
-For (5), the mean value inequality [10] implies that,
+The proof of (4) is the same as the proof of (3) in Theorem 4, we just omit it here.
+For (5), the mean value inequality [LS84] implies that,
 $M_{z_{ik}}(\frac{1}{2})=\sup\limits_{B(p_{k},\frac{1}{2})}|z_{ik}|\leq C(n),$
-holds for any $1\leq i\leq n$ and $k\in\mathbb{N}$. By the three-circle theorem in [13] and Cheng-Yau gradient estimate [5],
+holds for any $1\leq i\leq n$ and $k\in\mathbb{N}$. By the three-circle theorem in [Liu16c] and Cheng-Yau gradient estimate [CY75],
 $M_{z_{ik}}(r)\leq C(n)r^{d_{\operatorname{min}}^{(i)}},\quad|dz_{ik}|\leq C(n)r^{d_{\operatorname{min}}^{(i)}-1}$
 on $B_{p_{i}}(r)$ for any $1\leq i\leq n$, $k\in\mathbb{N}$ and $r\geq\frac{1}{2}$.
-Arzela-Ascoli theorem implies that $z_{ik}$ converges to a holomorphic function $z_{i\infty}$ of homogeneous degree no more than $d_{\operatorname{min}}^{(i)}$ uniformly on compact subsets.
+Arzela-Ascoli theorem implies that $z_{ik}$ converges to a holomorphic function $z_{i\infty}$ of homogeneous degree at most $d_{\operatorname{min}}^{(i)}$ uniformly on compact subsets.
 By (2.7) and the Gromov-Hausdorff convergence,
-$\min\limits_{\partial B(p_{\infty},r)}\sum\limits_{i=1}^{n}|z_{i\infty}|^{2}\geq c(n,\nu)r^{\frac{2}{D(n,\nu)}}>0.$ (2.5)
+(2.10) $\min\limits_{\partial B(p_{\infty},r)}\sum\limits_{i=1}^{n}|z_{i\infty}|^{2}\geq c(n,\nu)r^{\frac{2}{D(n,\nu)}}>0.$
 Firstly, we show these functions are algebraically independent. Suppose not, then there exists a nontrivial polynomial $P\in\mathbb{C}[x_{1},\cdots,x_{n}]$ such that
@@
 And $z_{i\infty}$ must of homogeneous degree $d_{\text{min}}^{(i)}$ by the dimension estimate in (3), because on $M_{\infty}$ the minimal degree of a n-tuple algebraically independent holomorphic homogeneous functions has to be $\overrightarrow{d_{\operatorname{min}}}$.
-Therefore, (4) implies that $(z_{1\infty},\cdots,z_{n\infty})$ must form a global coordinate as in (3).
-For (6), the first statement was proved in the proof of Theorem 3 in page 13 in [18].
-This concludes the proof of Conjecture 1.5 (2) in the maximal volume growth case.
+Therefore $(z_{1\infty},\cdots,z_{n\infty})$ must form a global coordinate as in (3).
+For (6), the first statement is proved in the proof of Theorem 3 in page 13 in [Liu24].
+This concludes the proof of Conjecture 2 (2) in the maximal volume growth case.
 The case of nonmaximal volume growth is trivial.
 ∎
-Remark 2.5.
-(1) As can be seen in the proof, if $M$ is of maximal volume growth, Conjecture 1.5 (2) still holds without the nonsplitting condition on the universal cover.
-(2) For simplicity, we just call the coordinate as in (2) in Theorem 2.4 a canonical coordinate on $M_{\infty}$, with degrees $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$.
+Remark.
+(1) As can be seen in the proof, if $M$ is of maximal volume growth, Conjecture 2 (2) still holds without the nonsplitting condition on the universal cover.
+(2) For simplicity, we just call the coordinate as in (2) in Theorem 5 a canonical coordinate on $M_{\infty}$, with degrees $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$.
 And we arrange them in non-decreasing order of their degrees.
 A natural corollary is the following homogeneity property of $\mathcal{O}_{P}(M)$:
-Corollary 2.6.
+Corollary 3.
 Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth. Then for any $f,g\in\mathcal{O}_{P}(M)$,
@@
 $\sup\limits_{B(p_{k},1)}|f_{ik}|=1.$
-Suppose $(M_{i},p_{i})\rightarrow(M_{\infty},p_{\infty})$ in the sense of Gromov-Hausdorff convergence, as in Proposition 5 in [15], each $f_{ik}\rightarrow f_{i\infty}$ with $f_{i\infty}$ is of homoegeneous degree $d_{\operatorname{min}}^{(i)}$.
+Suppose $(M_{i},p_{i})\rightarrow(M_{\infty},p_{\infty})$ in the sense of Gromov-Hausdorff convergence, as in Proposition 5 in [Liu18b], each $f_{ik}\rightarrow f_{i\infty}$ with $f_{i\infty}$ is of homoegeneous degree $d_{\operatorname{min}}^{(i)}$.
 As these functions generate $\mathcal{O}_{P}(M)$, we just need to prove: for any $f=f_{1}^{k_{1}}\cdots f_{n}^{k_{n}}$, $\operatorname{deg}(f)=\sum\limits_{i=1}^{n}k_{i}d_{\operatorname{min}}^{(i)}$.
@@
 $=a^{\sum\limits_{i=1}^{n}k_{i}d_{\operatorname{min}}^{(i)}}$
-Compare these we complete the proof. ∎
-Recently, Chu-Hao [7] obtain an optimal rigidity of the dimension estimate for polynomial growth holomorphic functions. Now we give an alternative proof, which seems much simpler.
-Theorem 2.7.
+Comparing these we complete the proof. ∎
+Recently, Chu-Hao [CH25] obtained an optimal rigidity of the dimension estimate for polynomial growth holomorphic functions. Now we give an alternative proof, which seems much simpler.
+Theorem 6 (Theorem 1.6 in [CH25]).
 Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature. Suppose there exists $d\geq 1$ such that
@@
 Suppose $M$ is of maximal volume growth, then take $(f_{1},\cdots,f_{n})$ be a canonical coordinate with minimal degree $(d^{(1)}_{\operatorname{min}},\cdots,d^{(n)}_{\operatorname{min}})$ with $d^{(i)}_{\operatorname{min}}\geq 1$ in non-decreasing order.
-From Corollary 2.6, $\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)=\#\{(m_{1},\cdots,m_{n})\in\mathbb{Z}^{n}_{\geq 0}\mid\sum\limits_{i=1}^{n}m_{i}d^{(i)}_{\min}\leq d\}$.
+From Corollary 3, $\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)=\#\{(m_{1},\cdots,m_{n})\in\mathbb{N}^{n}\mid\sum\limits_{i=1}^{n}m_{i}d^{(i)}_{\min}\leq d\}$.
 We know that $M$ is biholomorphically isometric to $\mathbb{C}^{n}$ if and only if $d^{(i)}_{\min}=1$ for all $i$.
@@
 Then $d^{(n)}_{\min}>1$, then
-$\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)$ $=\#\{(m_{1},\cdots,m_{n})\in\mathbb{Z}^{n}_{\geq 0}\mid\sum\limits_{i=1}^{n}m_{i}d^{(i)}_{\min}\leq d\}$
-$\leq\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(\mathbb{C}^{n})-\#\{(m_{1},\cdots,m_{n-1})\in\mathbb{Z}^{n-1}_{\geq 0}\mid\sum\limits_{i=1}^{n}m_{i}<d\}$
+$\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)$ $=\#\{(m_{1},\cdots,m_{n})\in\mathbb{N}^{n}\mid\sum\limits_{i=1}^{n}m_{i}d^{(i)}_{\min}\leq d\}$
+$\leq\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(\mathbb{C}^{n})-\#\{(m_{1},\cdots,m_{n-1})\in\mathbb{N}^{n-1}\mid\sum\limits_{i=1}^{n}m_{i}<d\}$
 $=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(\mathbb{C}^{n})-\binom{n+d-2}{d-1}$.
-with equality holds if and only $d^{(i)}_{\operatorname{min}}=1$ for $i\leq n-1$ and $1<d^{(n)}_{\operatorname{min}}<\frac{d}{d-1}$. This contradicts with our assumption.
+with equality holds if and only if $d^{(i)}_{\operatorname{min}}=1$ for $i\leq n-1$ and $1<d^{(n)}_{\operatorname{min}}<\frac{d}{d-1}$. This contradicts with our assumption.
 Now suppose the universal cover $\widetilde{M}$ splits.
-By the proof of Corollary 7.1 in [11], $\widetilde{M}=M_{1}^{m_{1}}\times M_{2}^{n-m_{1}}$, where $M_{1},M_{2}$ are both simply connected and has nonnegative bisectional curvature.
+By the proof of Corollary 7.1 in [Liu16a], $\widetilde{M}=M_{1}^{m_{1}}\times M_{2}^{n-m_{1}}$, where $M_{1},M_{2}$ are both simply connected and has nonnegative bisectional curvature.
 Also $M_{1}$ is of maximal volume growth.
@@
 And the $G$-action on $\mathbb{C}^{n}$ must be the deck transformation of $\pi_{1}(M)$.
-Claim: Let $\pi:\widetilde{M}\rightarrow M$ be the covering map, $f\in\mathcal{O}_{P}(M)$, then $\operatorname{deg}(f)=\operatorname{deg}(f\circ\pi)$.
-It’s obvious since $\pi(B(\tilde{p},r))=B(p,r)$ for any $\pi(\tilde{p})=p$ and any $r>0$.
+Claim 1.
+Let $\pi:\widetilde{M}\rightarrow M$ be the covering map, $f\in\mathcal{O}_{P}(M)$, then $\operatorname{deg}(f)=\operatorname{deg}(f\circ\pi)$.
+Proof of Claim 1. It’s obvious since $\pi(B(\tilde{p},r))=B(p,r)$ for any $\pi(\tilde{p})=p$ and any $r>0$. ∎
 Since $\mathcal{O}_{P}(M)$ is finitely generated, take a family of minimal generators of it, say $\{f_{1},\cdots,f_{N}\}\subset\mathcal{O}_{P}(\mathbb{C}^{n})$.
@@
 unless $\{f_{1},\cdots,f_{N}\}=A\{z_{1},\cdots,z_{n}\}+b$ for some constant $A,b$, in which case $\pi_{1}(M)$ is trivial. Therefore we complete the proof. ∎
-Remark 2.8.
-Similar argument can also prove the splitting theorem in Section 3 in [7], by using Theorem 4.1 in [23].
-2.2. Gradient expanding Kähler-Ricci solitons
-Definition 2.9.
+Remark.
+Similar argument can also prove the splitting theorem in Section 3 in [CH25], by using Theorem 4.1 in [NT03].
+2.2. Gradient expanding Kähler-Ricci solitons with nonnegative Ricci curvature
+Definition 3.
 A Kähler-Ricci soliton consists of a triple $(M,g,X)$, where $M$ is a Kähler manifold, $X$ is a holomorphic vector field on $M$, and $g$ is a complete Kähler metric on $M$ whose Kähler form $\omega$ satisfies
@@
 $\operatorname{Ric}_{i\bar{j}}-f_{i\bar{j}}+\lambda g_{i\bar{j}}=0,\quad f_{ij}=f_{\bar{i}\bar{j}}=0.$
-They we list some important properties of gradient expanding Kähler-Ricci solitons with nonnegative Ricci curvature which can be found in [2, 6, 30, 4, 28].
-Proposition 2.10.
+They we list some important properties of gradient expanding Kähler-Ricci solitons with nonnegative Ricci curvature which can be found in [CT05, Bry08, CCGGIIKLLN, Zhang09, CT11, Yang22].
+Proposition 1.
 Let $(N^{n},J,g,f)$ be a complete noncompact gradient expanding Kähler-Ricci solitons with nonnegative Ricci curvature. Then
@@
 (4) $\nabla^{g}f$ is a complete vector field. Let $\varphi(t)$ be a family of biholomorphism of $(N,J)$ with
-$\frac{\partial}{\partial t}\varphi(t-1)(x)=-\frac{1}{t}\nabla^{g}f(\varphi(t-1)(x)),\varphi(0)=\text{id}.$ (2.6)
+(2.11) $\frac{\partial}{\partial t}\varphi(t-1)(x)=-\frac{1}{t}\nabla^{g}f(\varphi(t-1)(x)),\varphi(0)=\text{id}.$
 It follows that
-$g(t)=t\varphi(t-1)^{*}g$ (2.7)
-solves the Kähler-Ricci flow equation on $[0,+\infty)$ with $g(1)=g$.
-(5) ((13) in [28]) The potential function $f$ satisfies the estimate
-$f(\varphi(O,t))\leq f(O)e^{2t}.$ (2.8)
+(2.12) $g(t)=t\varphi(t-1)^{*}g$
+solves the Kähler-Ricci flow equation on $0,+\infty)$ with $g(1)=g$.
+(5) ((13) in [Yang22]) The potential function $f$ satisfies the estimate
+(2.13) $f(\varphi(O,t))\leq f(O)e^{2t}.$
+(6) There exists a global coordinate called Poincaré coordinates on $N$ which is constructed as following: there exist local coordinates $\left(U_{O},z_{i}\right)$ at $O$ such that $\left\{\frac{\partial}{\partial z_{i}}\right\}$ is unitary at $O$ and the complexified vector field $\nabla f-\sqrt{-1}J\nabla f=\sum\limits_{i=1}^{n}\left(1+\mu_{i}\right)z_{i}\frac{\partial}{\partial z_{i}}$.
+Then one can extend $\left(U_{O},z_{i}\right)$ to the whole $(N,g)$ by defining
+$z_{i}(\psi(p,t))=e^{\left(1+\mu_{i}\right)t}z_{i}(p)$
+for any $p\in U_{O}$ where $\psi(p,t)$ is the holomorphic flow generated by $\nabla f$ starting from $p$.
 3. Proof of Theorem B
-The key observation is the following which may be well-known for experts in the fields. However, for reader’s convenience, we give a proof here.
-Lemma 3.1.
+The key observation is the following which may be well-known for experts in the field. However, for reader’s convenience, we give a proof here.
+Lemma 1.
 Let $(N,J,h,O,f)$ be a complete noncompact expanding Kähler-Ricci soliton with nonnegative bisectional curvature and we normalize it by assuming $f=R+|\nabla^{h(1)}f|^{2}$. Let
 $h(t)=t\varphi(t-1)^{*}h$
-be the unique Kähler-Ricci flow solution on $[0,+\infty)$ such that $h(1)=h$. Where
+be the unique Kähler-Ricci flow solution on $0,+\infty)$ such that $h(1)=h$. Where
 $\frac{\partial}{\partial t}\varphi(t-1)(x)=-\frac{1}{t}\nabla^{g}f(\varphi(t-1)(x)),\varphi(0)=\text{id}.$
@@
 Proof.
-Replace $(M,g)$ in Theorem 1.6 by $(N,h(1))$, we conclude that there exists a metric Kähler cone $N_{\infty}$ such that $(N,h(t),O)\rightarrow(N_{\infty},h_{\infty},O_{\infty})$ as $t\rightarrow 0$ in the Gromov-Hausdorff sense.
+Replace $(M,g)$ in Theorem 2 by $(N,h(1))$, we conclude that there exists a metric Kähler cone $N_{\infty}$ such that $(N,h(t),O)\rightarrow(N_{\infty},h_{\infty},O_{\infty})$ as $t\rightarrow 0$ in the Gromov-Hausdorff sense.
 So it remains to prove this cone $N_{\infty}$ is the unique tangent cone of $(N,h(1))$.
-In fact we have
-$(N_{\infty},h_{\infty},O_{\infty})$ $=\lim\limits_{t\to 0}(N,h(t),O)$
-$=\lim\limits_{t\to 0}(N,t\varphi(t-1)^{\star}h(1),O)$
-$=\lim\limits_{t\to 0}(N,th(1),\varphi(t-1)(O))$
-$=\lim\limits_{t\to 0}(N,th(1),(O)).$
-All the limits above are in the sense of Gromov-Hausdorff convergence. The second-to-last equality holds by the following claim:
 Claim 2.
 $\lim\limits_{t\to 0}d_{th(1)}(\varphi(t-1)(O),O)=0$.
-Proof.
-For any $s\leq 1$, by (2) and (2.8) in Proposition 2.10,
+Proof of Claim 2. For any $s\leq 1$, by (2) and (2.13) in Proposition 1,
 $|\nabla^{h(1)}f|^{2}\left(\varphi(s)(O)\right)\leq R+|f(\varphi(s)(O))|\leq C(1+e^{2})\leq C.$
@@
 $\sqrt{t}d_{h(1)}(\varphi(t-1)(O),O)\leq\sqrt{t}\int_{t}^{1}\frac{C}{s}ds=-C\sqrt{t}\log(t)\rightarrow 0\text{ as }t\rightarrow 0.$
-Thus, we now have a clear geometric framework characterizing the long-time asymptotics of the Kähler-Ricci flow on a complete noncompact Kähler manifold with bounded nonnegative bisectional curvature and maximal volume growth, along with its relationship to its tangent cones.
+In fact we have
+$(N_{\infty},h_{\infty},O_{\infty})$ $=\lim\limits_{t\to 0}(N,h(t),O)$
+$=\lim\limits_{t\to 0}(N,t\varphi(t-1)^{*}h(1),O)$
+$=\lim\limits_{t\to 0}(N,th(1),\varphi(t-1)(O))$
+$=\lim\limits_{t\to 0}(N,th(1),(O)).$
+All the limits above are in the sense of Gromov-Hausdorff convergence. The second-to-last equality holds by Claim 2: ∎
+Hence we now have a clear geometric framework characterizing the long-time asymptotics of the Kähler-Ricci flow on a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth, along with its relationship to its tangent cones.
 Now we prove Theorem B.
-Proof of Theorem B:
-Following the notation of Theorem 1.6, take the Poincaré coordinate $(w_{1},\cdots,w_{n})$ as in page 2622 in [28] on $(N,h(1),O)$ and normalize them such that
-$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B_{h(1)}(O,1)}{w_{i}}\overline{w_{j}}=\delta_{ij},\quad w_{i}(O)=0\quad\text{ for any }i,j\in\{1,\cdots,n\}.$
-According to the proof of Theorem 1.4 in [28], the aforementioned Poincaré coordinate $(w_{1},\cdots,w_{n})$ is a canonical coordinate on $(N,h(1),O)$.
-Following the procedure of (5) in Theorem 2.4 on $(N,h(1),O)$ and $(M,g,p)$, we get two canonical coordinates $(z_{1\infty},\cdots,z_{n\infty})$ and $(w_{1\infty},\cdots,w_{n\infty})$ on $N_{\infty}=M_{\infty}$, with degree $\overrightarrow{d_{\operatorname{min}}}(M)$ and $\overrightarrow{d_{\operatorname{min}}}(N)$.
-Thus from (4) in Theorem 2.4,
+Proof of Theorem B.
+Following the notation of Theorem 2, take the Poincaré coordinate $(w_{1},\cdots,w_{n})$ from Proposition 1 on $(N,h(1),O)$ and normalize them such that
+$\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B_{h(1)}(O,1)}{w_{i}}\overline{w_{j}}=\delta_{ij},\quad w_{i}(O)=0\quad\text{ for any }i,j\in\{1,\cdots,n\}.$
+According to the proof of Theorem 1.4 in [Yang22], the aforementioned Poincaré coordinate $(w_{1},\cdots,w_{n})$ is a canonical coordinate on $(N,h(1),O)$.
+Following the procedure of (5) in Theorem 5 on $(N,h(1),O)$ and $(M,g,p)$, we get two canonical coordinates $(z_{1\infty},\cdots,z_{n\infty})$ and $(w_{1\infty},\cdots,w_{n\infty})$ on $N_{\infty}=M_{\infty}$, with degree $\overrightarrow{d_{\operatorname{min}}}(M)$ and $\overrightarrow{d_{\operatorname{min}}}(N)$.
+Thus from (4) in Theorem 5,
 $\overrightarrow{d_{\operatorname{min}}}(M)=\overrightarrow{d_{\operatorname{min}}}(N)=\{\mu_{1}+1,\cdots,\mu_{n}+1\}.$
 Therefore, (1),(2) has been proved.
-For (3), note that in [18], Liu has proved that $\operatorname{ASCD}(M,g(t))=4n(\sum\limits_{i=1}^{n}d^{(i)}_{\operatorname{min}}-n)=4n(\sum\limits_{i=1}^{n}\mu_{i})$ is invariant along $t$.
+For (3), note that in [Liu24], Liu has proved that $\operatorname{ASCD}(M,g(t))=4n(\sum\limits_{i=1}^{n}d^{(i)}_{\operatorname{min}}-n)=4n(\sum\limits_{i=1}^{n}\mu_{i})$ is invariant along $t$.
 The proof is complete.
-As an application, we can use it to prove Corollary 1.11:
-Proof of Corollary 1.11:
-Firstly, we assume $(M,g)$ has bounded curvature.
+∎
+As an application, we can use it to prove Corollary 1:
+Proof of Corollary 1:
 Take a canonical coordinate $(f_{1},\cdots,f_{n})$ on $(M,g)$, with degree $\overrightarrow{d_{\operatorname{min}}}$.
-By comparing the distance function $d_{g(t)}(x,p)$ along the flow using the shrinking balls lemma by Simon and Topping (Corollary 3.3 in [26]), we may conclude that $(f_{1},\cdots,f_{n})$ is still a global coordinate on $(M,g(t))$, with the same degree $\overrightarrow{d_{\operatorname{min}}}$.
-Let $g_{k}(t)=\frac{1}{t_{i}}g(t_{i}t)$, we normalized them such that
-$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B_{g_{k}(1)}(O,1)}{f_{ik}}\overline{f_{jk}}=\delta_{ij},\quad f_{ik}(p)=0\quad\text{ for any }i,j\in\{1,\cdots,n\},k\in\mathbb{N}.$
+By comparing the distance function $d_{g(t)}(x,p)$ along the flow using the shrinking balls lemma by Simon and Topping (Corollary 3.3 in [ST22]), we may conclude that $(f_{1},\cdots,f_{n})$ is still a global coordinate on $(M,g(t))$, with the same degree $\overrightarrow{d_{\operatorname{min}}}$.
+Let $g_{k}(t)=\frac{1}{t_{i}}g(t_{i}t)$, we normalize them such that
+$\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B_{g_{k}(1)}(O,1)}{f_{ik}}\overline{f_{jk}}=\delta_{ij},\quad f_{ik}(p)=0\quad\text{ for any }i,j\in\{1,\cdots,n\},k\in\mathbb{N}.$
 which makes $(f_{1k},\cdots,f_{nk})$ serving as a canonical coordinate on $(M,g_{i}(1))$.
-Following the same proof as in (5) in Theorem 2.4, these $(f_{1k},\cdots,f_{nk})$ converges smoothly to $(f_{1\infty},\cdots,f_{n\infty})$ on $(N,h(1))$ and serving as a canonical coordinate on $N$.
-Due to (3) in Theorem 2.1, WLOG we can assume $(f_{1\infty},\cdots,f_{n\infty})$ is just a Poincaré coordinate on $(N,h(1))$.
+Following the similar proof as in (5) in Theorem 5, these $(f_{1k},\cdots,f_{nk})$ converges smoothly to $(f_{1\infty},\cdots,f_{n\infty})$ on $(N,h(1))$ and serves as a canonical coordinate on $N$.
+Due to (3) in Theorem 4, we can assume $(f_{1\infty},\cdots,f_{n\infty})$ is just a Poincaré coordinate on $(N,h(1))$.
 Therefore
 $\operatorname{Ric}_{h(1)}\left(\frac{\frac{\partial}{\partial f_{i\infty}}}{|\frac{\partial}{\partial f_{i\infty}}|},\overline{\frac{\frac{\partial}{\partial f_{i\infty}}}{|\frac{\partial}{\partial f_{i\infty}}|}}\right)=\mu_{i}.$
-Finally, by the smooth convergence of $g_{i}(1)$ and the proof of Theorem 1.2 in [3], the Corollary is proved.
+Finally, by the smooth convergence of $g_{i}(1)$ and the proof of Theorem 1.2 in [CT06], the Corollary is proved.
 In fact, for any $i$ and any nonzero $v\in V_{1}\oplus\cdots\oplus V_{i}$ but $v\notin V_{1}\oplus\cdots\oplus V_{i-1}$,
 $\lim\limits_{t\to\infty}t\operatorname{Ric}(v(t),\overline{v(t)})=\mu_{i}$
-always holds.
-If $(M,g)$ has unbounded curvature, thanks to Theorem 1.9, above argument can still be applied.
+always holds. ∎
 4. Proof of Theorem A
 Proof of Theorem A:
-Firstly we assume $M$ is of nonmaximal volume growth, assume there exists a nonzero polynomial holomorphic $n$-form $s\in H^{0}_{P}(M,\mathcal{K}_{M})$, the Poincaré-Lelong equation states that
+Firstly we assume $M$ is of nonmaximal volume growth, assume there exists a nonzero polynomial growth holomorphic $n$-form $s\in H^{0}_{P}(M,\mathcal{K}_{M})$, the Poincaré-Lelong equation states that
 $\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\log\|s\|^{2}=(s)+\operatorname{Ric}\geq 0,$
 where $(s)$ is the zero divisor of $s$.
-In other words, $h=\log(\|s\|^{2}+1)$ is a smooth plurisubharmonic function of logarithmic growth.
-Since the universal cover of $M$ does not split, the heat-flow method of Ni-Tam [23] can be applied to perturb $h$ to $h^{\prime}$ which is a smooth plurisubharmonic function of still logarithmic growth (for details, see Theorem 0.1 (iii) in [23]).
-Then we solve the $\bar{\partial}$-equation by applying the Hörmander $L^{2}$-estimate similarly as the proof of Lemma 3 in [12], there exists a nonconstant polynomial growth holomorphic function on $M$, contradicting with Theorem 2 in [12].
+In other words, $h:=\log(\|s\|^{2}+1)$ is a smooth plurisubharmonic function of logarithmic growth.
+Since the universal cover of $M$ does not split, the heat-flow method of Ni-Tam [NT03] can be applied to perturb $h$ to $h^{\prime}$ which is a smooth plurisubharmonic function of still logarithmic growth.
+Then we solve the $\bar{\partial}$-equation by applying the Hörmander $L^{2}$-estimate to obtain a nonconstant polynomial growth holomorphic function on $M$, contradicting with Theorem 1.
 As a result, $D_{\operatorname{min}}=+\infty=\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n$.
 Now we assume $M$ is of maximal volume growth.
-Let $(z_{1},\cdots,z_{n})$ be a canonical coordinate of $M$ with degrees $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$ as stated in Remark 2.3.
-By gradient estimate [5],
+Let $(z_{1},\cdots,z_{n})$ be a canonical coordinate of $M$ with degrees $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$.
+By gradient estimate [CY75],
 $\operatorname{deg}(dz_{1}\wedge\cdots\wedge dz_{n})\leq\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n,$
 therefore $D_{\operatorname{min}}\leq\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n$.
-The proof of the opposite direction is similar to the proof of Corollary 1.5 in [28], but a little more complicated.
-Let $s\in H^{0}_{D_{\operatorname{min}}}(M,\mathcal{K}_{M})$, then note that Corollary 3 in [13] also holds for holomorphic sections of holomorphic line bundle with nonpoaitive curvature, there exists a constant $C>0$ such that
-$M_{s}(r,0)\leq Cr^{D_{\operatorname{min}}}\quad\text{ for any }r>0.$ (4.1)
-Fix $t_{k}\rightarrow+\infty$, there exists a long-time Kähler-Ricci flow solution $g(t)$ which satisfies properties in Theorem 1.9. And set $g_{k}(t)=\frac{1}{t_{i}}g(t_{i}t)$, assume $t_{k}\geq 1$.
+We will give two distinct proofs of the opposite direction:
+Proof I.
+The first proof of the opposite direction is similar to the proof of Corollary 1.5 in [Yang22], but a little more complicated.
+Let $s\in P(M,\mathcal{K}_{M})$ with degree $D_{\operatorname{min}}$, then note that Corollary 3 in [Liu16c] can also apply on holomorphic sections of holomorphic line bundle with nonpositive curvature, there exists a constant $C>0$ such that
+(4.14) $M_{s}(r,0)\leq Cr^{D_{\operatorname{min}}}\quad\text{ for any }r>0.$
+Fix $t_{k}\rightarrow+\infty$, there exists a long-time Kähler-Ricci flow solution $g(t)$ which satisfies properties in Theorem 2. And set $g_{k}(t)=\frac{1}{t_{i}}g(t_{i}t)$, assume $t_{k}\geq 1$.
 Claim 3.
 […]
-Proof.
-By the shrinking balls lemma by Simon and Topping (Corollary 3.3 in [26]), for any $t_{k}>0$, we have
+Proof of Claim 3.
+By the shrinking balls lemma by Simon and Topping (Corollary 3.3 in [ST22]), for any $t_{k}>0$, we have
 $d_{g}\leq d_{g(t_{k})}+C^{\prime}_{k}\sqrt{t_{k}}$
 on $M\times[0,+\infty)$.
-Also, thank to Corollary 1.10, (1.2) also holds for $M$ with unbounded curvature, which means that
+Then by (1.2) we obtain that
 $\lim\limits_{t\to\infty}\log\frac{\det(g_{i\bar{j}}(x,t))}{\det(g_{i\bar{j}}(x,0))}=-C.$
-for some nonnegative constant $C$ which relies only on $M$ and is independent of $x$. Also by the Ricci flow equation,
+for some nonnegative constant $C$ which relies only on $M$ and is independent of $x$. From Ricci flow equation,
 $\frac{d}{dt}\log\frac{\det(g_{i\bar{j}}(x,t))}{\det(g_{i\bar{j}}(x,0))}=-R(x,t)\leq 0.$
@@
 The claim is proved. ∎
-[…]
-Thanks to Theorem 1.9, the boundedness condition of curvature in Theorem 1.6 can be removed.
-Using the notations in Theorem 1.6, recall the sequence $(U_{k},F_{k}^{*}J.F_{k}^{*}(g_{k}(1)),O)$ converges smoothly to $(N,J_{\infty},h(1),O)$ uniformly on compact subsets.
-Define
-$s_{k}=\frac{s\mid_{F_{k}(U_{k})}}{\sup\limits_{B_{g_{k}(1)}(p,1)}\|s\|},\quad\tilde{s}_{k}=F_{k}^{*}s_{k}.$ (4.2)
-Since $\mathcal{K}_{M}(g_{k}(1))$ has nonpositive curvature, three-circle theorem in [13] can be applied to get for any fixed $r>0$, for large $k$,
+Using the notations in Theorem 2, recall the sequence $(U_{k},F_{k}^{*}J.F_{k}^{*}(g_{k}(1)),O)$ converges smoothly to $(N,J_{\infty},h(1),O)$ uniformly on compact subsets. Define
+(4.15) $s_{k}=\frac{s\mid_{F_{k}(U_{k})}}{\sup\limits_{B_{g_{k}(1)}(p,1)}\|s\|},\quad\tilde{s}_{k}=F_{k}^{*}s_{k}.$
+Since $\mathcal{K}_{M}(g_{k}(1))$ has nonpositive curvature, three-circle theorem can be applied to get for any fixed $r>0$, for large $k$,
 $\sup\limits_{B_{g_{k}(1)}(p,r)}\|s_{k}\|\leq r^{D_{\operatorname{min}}}.$
 After pulling back by $F_{k}$, we obtain
-$\sup\limits_{B_{F_{k}^{*}g_{k}(1)}(O,r)}\|s_{k}\|\leq r^{D_{\operatorname{min}}}.$ (4.3)
+(4.16) $\sup\limits_{B_{F_{k}^{*}g_{k}(1)}(O,r)}\|s_{k}\|\leq r^{D_{\operatorname{min}}}.$
 holds for large $k$.
 Fix a nonvanishing holomorphic $n$-form $\Omega$ on $N$, then write $s_{k}=f_{k}\Omega_{k}$ on $U_{k}$, where $f_{k}$ is a holomorphic function with respect to $F_{k}^{*}J$ on $U_{k}$, $\Omega_{k}$ is a nonvanishing holomorphic $n$-form with respect to $F_{k}^{*}J$ and converges to $\Omega$ uniformly on each fixed compact subset in $N$ (For example, take $U_{k}$ be a sequence of exhausted Euclidean balls in $N\cong\mathbb{C}^{n}$).
-Therefore there exists a constant $C=C(K)$ such that $\|\Omega_{k}\|_{F_{k}^{*}(g_{k}(1))}\geq C(K)$ uniformly on $K\subset N$ compact. Then we can deduces
+Therefore there exists a constant $C=C(K)$ such that $\|\Omega_{k}\|_{F_{k}^{*}(g_{k}(1))}\geq C(K)$ uniformly on $K\subset N$ compact. Then we can deduce
 $\sup\limits_{K}|f_{k}|\leq\frac{1}{C(K)}r^{D_{\operatorname{min}}}$
@@
 $D_{\operatorname{min}}=\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n.$
-A much simpler way is consider directly a tangent cone of $M$.
+Proof II.
+A much simpler way is considering directly a tangent cone of $M$.
 Assume $(M,g_{i}=r_{i}^{-2}g)\rightarrow M_{\infty}$ in the Gromov-Hausdorff sense, where $r_{i}\rightarrow\infty$.
@@
 $D_{\operatorname{min}}=\operatorname{deg}(dz_{1}\wedge\cdots\wedge dz_{n}).$
-And from (4) we have known that
+We have known that
 $D_{\operatorname{min}}\leq\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n.$
-By the three-circle theorem, assume
+Assume
 $D_{\operatorname{min}}\leq\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n-\varepsilon$
 for some $\varepsilon>0$.
-Following the notations in Theorem 2.3 (5), and define
+Following the notations in Theorem 5 (5), and define
 $\varphi_{i}=\log\|dz_{1i}\wedge\cdots\wedge dz_{ni}\|^{2}_{g_{i}}$
-As in [18], since $\varphi_{i}$ are psh on $M_{i}$, by passing to a subsequence, we may assume $\varphi_{i}$ converges to a psh function on $M_{\infty}$ in $L_{\operatorname{loc}}^{1}$, denoted by $\varphi=\log|dz|^{2}$.
+As in [Liu24], since $\varphi_{i}$ is plurisubharmonic on $M_{i}$, by passing to a subsequence, we may assume $\varphi_{i}$ converges to a plurisubharmonic function on $M_{\infty}$ in $L_{\operatorname{loc}}^{1}$, denoted by $\varphi=\log|dz|^{2}$.
 Therefore $e^{\varphi_{i}}\rightarrow e^{\varphi}$ almost everywhere.
-By the three circle theorem as in the proof of Theorem 2.3 (5), there exists a $C$ such that
+By the three circle theorem as in the proof of Theorem 5 (5), there exists a $C$ such that
 $e^{\varphi_{i}}\leq Cr_{i}^{2D_{\operatorname{min}}}\leq Cr_{i}^{2(\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n-\varepsilon)}.$
 uniformly.
-Thus $e^{\varphi}\leq Cr^{2(\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n-\varepsilon)}$ almost everywhere, which contradicts with Claim 1 in [18] if $\varepsilon>0$, since $e^{\varphi}$ is homogeneous of degree $2(\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n)$.
+Thus $e^{\varphi}\leq Cr^{2(\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n-\varepsilon)}$ almost everywhere, which contradicts with Claim 1 in [Liu24] if $\varepsilon>0$, since $e^{\varphi}$ is homogeneous of degree $2(\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n)$.
 This directly implies that $D_{\operatorname{min}}=\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n.$
+5. Tangent cones and further questions
+In this section, we will try to study the structure of the tangent cones of a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth.
+In [Kap02], Kapovitch proved that, Let $M_{\infty}=C(X)$ be a tangent cone of a complete noncompact Riemannian manifold with nonnegative sectional curvature and maximal volume growth, then the link $X$ has Alexandrov curvature bounded below by 1 and is homeomorphic to a sphere.
+Firstly, we will prove Theorem C, which is a Kähler analogue of Kapovitch’s result.
+Proof of Theorem C (1).
+To begin with, we recall the definition of so called weighted Sasaki spheres, see [HS16] for a precise description.
+Let $\mathbb{C}^{n}$ equips with the flat metric $\omega_{0}=\frac{\sqrt{-1}}{2}\sum\limits_{i=1}^{n}dz^{i}\wedge d\bar{z}^{i}$.
+The link, standard sphere $\mathbb{S}^{2n-1}$, is a Sasaki manifold with Reeb vector field $\xi_{0}=\sum\limits_{i=1}^{n}(y^{i}\frac{\partial}{\partial x^{i}}-x^{i}\frac{\partial}{\partial y^{i}})=\operatorname{Re}(\sqrt{-1}\sum\limits_{i=1}^{n}z^{i}\frac{\partial}{\partial z_{i}})$ and contact form $\eta_{0}=\frac{1}{\sum\limits_{i=1}^{n}|z_{i}|^{2}}\sum\limits_{i=1}^{n}(y^{i}dx^{i}-x^{i}dy^{i})$.
+The Reeb cone in $\mathrm{t}^{*}=\mathbb{R}^{n}$ in this case is just $\mathcal{R}=\mathbb{R}_{+}^{n}$.
+For any $(d_{1},\cdots,d_{n})\in\mathbb{R}_{+}^{n}$, apply Type I deformation (see Section 2 in [HS16]) to get a new Kähler cone metric $\omega^{\prime}=\frac{\sqrt{-1}}{2}\partial\bar{\partial}{r^{\prime}}^{2}$ on $\mathbb{C}^{n}$ with Reeb field $\xi^{\prime}=\sum\limits_{i=1}^{n}d_{i}(y^{i}\frac{\partial}{\partial x^{i}}-x^{i}\frac{\partial}{\partial y^{i}})=\operatorname{Re}(\sqrt{-1}\sum\limits_{i=1}^{n}d_{i}z^{i}\frac{\partial}{\partial z_{i}})$.
+The new link $\{r^{\prime}=1\}$ is still a Sasaki manifold and diffeomorphic to $\mathbb{S}^{2n-1}$ which can be seen in the proof of Lemma 2.2 in [HS16].
+We will call this Sasaki metric on $\mathbb{S}^{2n-1}$ a standard weighted sphere metric.
+Now we are going to prove Theorem C (1).
+Let $(M_{\infty}=C(X)\cong\mathbb{C}^{n},\omega=\frac{\sqrt{-1}}{2}\partial\bar{\partial}r^{2})$ be a tangent cone of $M$ with $\omega$ be the Kähler current.
+There exists a homogeneous coordinate $(z_{1},\cdots,z_{n})$ on $M_{\infty}$ with homogeneous degree $(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)})$ and the Reeb field $\xi=\operatorname{Re}(\sqrt{-1}\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}z^{i}\frac{\partial}{\partial z_{i}})$.
+Then replace $(d_{1},\cdots,d_{n})$ in the last paragraph by $(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)})$ to get a smooth Kähler cone metric $\omega^{\prime}=\frac{\sqrt{-1}}{2}\partial\bar{\partial}{r^{\prime}}^{2}$ over a standard weighted sphere with the same Reeb field $\xi$.
+Denote the flows generated by $r\frac{\partial}{\partial r}$ and $\xi=Jr\frac{\partial}{\partial r}$ by $\Phi$ and $\Psi$.
+Note that they are both independent of these two cone metrics.
+[…]
+In particular, the link $X=\{r=1\}=\{r^{\prime}=\exp(-\phi)\}$ is homeomorphic to the link $\{r^{\prime}=1\}$, which is a standard weighted sphere thus diffeomorphic to $\mathbb{S}^{2n-1}$. ∎
+If a smooth Kähler cone has nonnegative bisectional curvature, a simple computation implies that the transverse bisectional curvature $\operatorname{BK}^{T}\geq 2$.
+The main theorem in [HS16] says that
+Theorem 7 (Theorem 1.1 in [HS16]).
+Let $(M^{2n-1},g)$ be a compact simply connected Sasaki manifold with positive bisectional curvature, then its Kähler cone is biholomorphic to $\mathbb{C}^{n}$.
+Moreover $M$ is a weighted Sasaki sphere, which is obtained through a type II deformation from the standard weighted Sasaki sphere defined above.
+The following is defined by Van Coevering in [Van15] to develop pluripotential theory on Sasaki manifolds.
+Definition 4.
+Let $M$ be a Sasaki manifold, $\theta$ be a closed basic (1,1) form.
+A function $u\colon M\rightarrow\mathbb{R}\cup\{-\infty\}$ is said to be transversally $\theta$-plurisubharmonic ($\theta$-psh) if $u$ is invariant under the Reeb flow, is upper semicontinuous, in each foliation chart is quasi-plurisubharmonic and
+$\theta+d_{B}d_{B}^{c}u\geq 0,$
+as a $(1,1)$-current. The set of $\theta$-psh functions on $M$ is denoted by $\operatorname{PSH}(M,\theta)$.
+Using the notations in the proof of Theorem C (1), on the Kähler cone $C(S)$ over the standard weighted Sasaki sphere $S$ with Reeb field $\xi=\operatorname{Re}(\sqrt{-1}\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}z^{i}\frac{\partial}{\partial z_{i}})$, contact form $\eta^{\prime}=d^{c}\log r^{\prime}$ and Kähler metric $\omega^{\prime}=\frac{\sqrt{-1}}{2}\partial\bar{\partial}{r^{\prime}}^{2}$.
+The transverse Kähler form ${\omega^{\prime}}^{T}=\frac{1}{2}d\eta^{\prime}=\sqrt{-1}\partial\bar{\partial}\log r^{\prime}$.
+According to Proposition 1 in [Liu24], the current $\sqrt{-1}\partial\bar{\partial}\log r=\sqrt{-1}\partial\bar{\partial}\log(r^{\prime}\exp\phi)={\omega^{\prime}}^{T}+\frac{1}{2}d_{B}d_{B}^{c}\phi\geq 0$, which means that the basic function $\phi\in\operatorname{PSH}(S,{\omega^{\prime}}^{T})$.
+In some sense, we can regard the link $X$ over the tangent cone $M_{\infty}=C(X)$ as a singular weighted Sasaki sphere.
+So it’s natural to ask that
+Problem 1.
+Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth.
+Let $(M_{\infty}\cong\mathbb{C}^{n},\omega=\frac{\sqrt{-1}}{2}\partial\bar{\partial}r^{2},d)$ be a tangent cone at infinity of $M$ with Reeb field $\xi$.
+If there exists a sequence of smooth Kähler cone metrics $(\mathbb{C}^{n},\omega_{i}=\frac{\sqrt{-1}}{2}\partial\bar{\partial}r_{i}^{2})$, which have the same Reeb field with $M_{\infty}$ such that $\operatorname{BK}(\omega_{i})\geq-\frac{\Psi(\frac{1}{i})}{r_{i}^{2}}$ and $(\mathbb{C}^{n},\omega_{i})$ converges to $(M_{\infty},d)$ in the Gromov-Hausdorff sense?
+The main difficulty here is the insufficient regularity of the Kähler potential $r^{2}$ on $M_{\infty}$.
+In general, one needs at least $C^{2}$ regularity of the limit metric in order to define curvature in the classical sense.
+However, thanks to Lott’s work (Definition 2 (3) and Theorem 5 (1)), it suffices to establish a pluripotential–theoretic version of the condition $\operatorname{BK}\geq K$”.
+This allows us to reduce the regularity requirement of the limit metric.
+Moreover, by Anderson [And89], the assumption of quadratic curvature decay guarantees exactly this level of regularity.
+However, it seems still hard to get uniform lower bound for bisectional curvature.
+Proof of Theorem C (2). We follow the notations in the proof of Theorem C (1).
+Since the curvature is of quadratic decay, the limit metric has $C^{1,\alpha}$ regularity.
+So the Kähler potential $\frac{r^{2}}{2}$, thus the basic function $\phi$ has $C^{3,\alpha}$ regularity.
+Following the proof of Proposition 2.2.6 in [Van15], there exists a sequence of basic smooth functions $\phi_{j}\in\operatorname{PSH}(S,{\omega^{\prime}}^{T})$ decreasing to $\phi$.
+In fact, we can require that $\phi_{j}\rightarrow\phi$ in $C^{3}$-topology.
+This is because when following the proof of Proposition 2.2.6 in [Van15], when using the regularization process described in Blocki-Kolodziej [BK07], since we work on a precompact domain in $M_{\infty}\cong\mathbb{C}^{n}$, there is no need to use the “cut-off function” and “regularized maximum”, as these might disrupt the convergence of the derivatives.
+the standard smooth regularization by convolution will suffice.
+So in particular, for any $p\in\mathbb{C}^{n}$, the distance functions $d^{i}_{p}\rightarrow d$ uniformly on compact subsets.
+∎
+Now we know that given a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth, any tangent cone of it is a (not necessarily smooth) Kähler cone metric on $\mathbb{C}^{n}$ with nonnegative bisectional curvature.
+Conversely, one may ask that for which Kähler cone metric on $\mathbb{C}^{n}$ is it possible to find such a manifold with such cone can be realized as a tangent cone of this manifold?
+For simplicity we consider smooth Kähler cones.
+By Theorem 5 (2), it’s equivalent to ask,
+Problem 2.
+For any smooth Kähler cone metric $(\mathbb{C}^{n},\omega=\frac{\sqrt{-1}}{2}\partial\bar{\partial}r^{2})$ with nonnegative bisectional curvature, can we find a gradient expanding Kähler-Ricci soliton with nonnegative bisectional curvature coming out of the cone?
+The problem can be seen as a Kähler analogue of a result of Deruelle [Der16], which smooths Riemannian cone with positive curvature operator out by expanding Ricci solitons.
+When the Kähler cone is regular, and has positive curvature operators on (1,1)-forms, above problem has been solved by Conlon-Deruelle in [CD20] by solving a smooth family of Monge-Ampére equations using continuity method.
+For more related work, we refer to [CDS24].